### v2

# On the distance problem over finite p-adic rings

Abstract
In this paper, we study the distance problem in the setting of finite p-adic rings. In odd dimensions, our results are essentially sharp. In even dimensions, we clarify the conjecture and provide examples to support it. Surprisingly, compared to the finite field case, in this setting, we are able to provide a large family of sets such that the distance conjecture holds. By developing new restriction type estimates associated to circles and orbits, with a group theoretic argument, we will prove the $4/3$-parallel result in the two dimensions. This answers a question raised by Alex Iosevich. In a more general scenario, the existence/distribution of geometric/graph configurations will be also considered in this paper. The main results present improvements and extensions of the recent work due to Ben Lichtin (2019, 2023). In comparison with Lichtin’s method, our approach is much simpler and flexible, which is also one of the novelties in this paper.
1 Introduction
Let $p$ be a prime, $r$ be a positive integer, and $\mathbb{Z}/p^{r}\mathbb{Z}$ be a finite $p$-adic ring. For a postive integer $n$, the “distance” between two points $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, denoted by $||\mathbf{x}-\mathbf{y}||$, is defined by
$||\mathbf{x}-\mathbf{y}||=(x_{1}-y_{1})^{2}+\cdots+(x_{n}-y_{n})^{2}\quad(\textsf{mod }p^{r}).$
Given subsets $E_{1},E_{2}\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, the distance set determined by points in $E_{1}\times E_{2}$ is denoted by $\Delta_{n,r}(E_{1},E_{2})$, i.e.,
$\Delta_{n,r}(E_{1},E_{2})=\{\|\mathbf{x}-\mathbf{y}\|:\,\mathbf{x}\in E_{1},\,\mathbf{y}\in E_{2}\}.$
For simplicity, we write $\Delta_{n,r}(E)=\Delta_{n,r}(E,E)$. We also denote the density by $\delta_{E_{1},E_{2}}:=\frac{\sqrt{|E_{1}||E_{2}|}}{p^{rn}}$ and $\delta_{E}:=\frac{|E|}{p^{rn}}$.
In the setting of finite $p$-adic rings, the Erdős-Falconer distance problem is stated as follows.
Question 1.1.
What is the smallest density threshold $\delta\in(0,1)$ independent of $r$ such that $|\Delta_{n,r}(E)|\gg p^{r}$ whenever $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq\delta$?
This problem was initially studied in the case $r=1$, i.e. over finite fields, due to Iosevich and Rudnev [7]. More precisely, they proved that if $|E|\geq Cp^{\frac{n+1}{2}}$ for some sufficient large constant $C$, then the distance set $\Delta_{n,1}(E)$ covers the whole field. Hart, Iosevich, Koh, and Rudnev [4] indicated that the exponent $\frac{n+1}{2}$ is sharp in odd dimensions. In even dimensions, it is conjectured that the right exponent should be $n/2$. In two dimensions, Chapman, Erdogan, Hart, Iosevich, and Koh [2] proved the exponent $4/3$ by using an extension theorem associated to circles in the plane. This result was recently improved to $5/4$ by Murphy, Petridis, the first listed author, Rudnev, and Stevens in [20] by using algebraic methods and results from incidence geometry.
When $r>1$, by extending the techniques from finite fields, Covert, Iosevich and Pakianathan [1] proved that $U_{r}\subset\Delta_{n,r}(E)$ whenever $\delta_{E}\gg r(r+1)p^{-\frac{n-1}{2}}$, where $U_{r}:=(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$ is the set of units. This result is only non-trivial when $r$ is bounded, and does not offer a uniform density independent of $r$, to Question 1.1. By using a different and sophisticated approach, namely, a combination of $p$-adic analysis and estimates for a class of exponential sums mod $p$, Lichtin [15] proved that $U_{r}\subset\Delta_{n,r}(E)$ if $\delta_{E}\gg p^{-\frac{n-1}{2}}$. As mentioned in his paper, the main advantage of his approach is that the argument detects nontrivial cancellations within certain exponential sums mod $p^{r}$ which were not used in the work of Covert et al. [1].
In this paper, the first purpose is to study the distance problem in a general setting when the distance function is replaced by a much more general polynomial. In the case of the usual distance function, one of the novelties of this paper is to present a much simpler and flexible approach than that of Lichtin. Our results are essentially sharp in odd dimensions. The second purpose of this paper is to clarify the conjecture in even dimensions and provide examples to support it. In two dimensions, by developing new restriction type estimates associated to circles and orbits, with a group theoretic argument, we will prove the $4/3$-parallel result. Surprisingly, in comparison with the finite field case, in this setting, we are able to provide a large family of sets such that the distance conjecture holds. The third purpose is to study the existence/distribution of geometric/graph configurations. To be precise, we are interested in the following questions.
Question 1.2.
Let $H$ be a given graph, $E$ be any subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ and $j$ be any element of $U_{r}$. What is the smallest density threshold $\delta\in(0,1)$ independent of $r$ such that $E$ contains a copy of $H$ at distance $j$ whenever $\delta_{E}\gg\delta$?
Question 1.3.
Let $\mathcal{C}$ be a given geometric configuration and $E$ be any set in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$. What is the smallest density threshold $\delta\in(0,1)$ not depending on $r$ such that $E$ contains a copy of $\mathcal{C}$ whenever $\delta_{E}\gg\delta$?
This paper addresses the graphs of being cycles, chains, and trees, and a geometric configuration of rectangles. In contrast to initial methods/results over finite fields, the main challenge one has to deal with in this setting is to find an effective approach to obtain a uniform density which does not depend on $r$. Lichtin’s method in [15] is based on a combination of $p$-adic analysis and estimates for a class of exponential sums mod $p$. To find cancellations in exponential sums, he investigated paths of points (descending family of neighbourhoods), and his arguments are heavily relied on computing Hessian matrix of all levels. The approach we introduce in this paper is much simpler, which involves polynomial congruences and Fourier analysis in finite rings. Our arguments only rely on the first order Jacobian matrix, and avoid the second order Hessian matrix.
1.1 Results on the generalized distance sets
For a polynomial $F(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ in $n$ variables. We denote $F_{i}(\mathbf{x})=\frac{\partial F}{\partial x_{i}}(\mathbf{x})$ for all $1\leq i\leq n$, and denote $(\nabla F)(\mathbf{x})=(F_{1}(\mathbf{x}),F_{2}(\mathbf{x}),\ldots,F_{n}(\mathbf{x}))$. Our first result reads as follows.
Theorem 1.4.
Let $F(\mathbf{x})$ be a given polynomial in $\mathbb{Z}[\mathbf{x}]$ in $n\geq 2$ variables. Suppose that the following conditions hold for some positive constants $c_{1},c_{2},c_{3}$ with $c_{1}<1$, for some prime $p$, and for some $j\in\mathbb{Z}$ with $(j,p)=1$:
(i) $(\nabla F)(\mathbf{x})\not\equiv\mathbf{0}\,(\textsf{mod }p)$ when $F(\mathbf{x})\equiv j\,(\textsf{mod }p)$;
(ii) |#$\big{|}\#\{\mathbf{x}\,(\textsf{mod }p):\,F(\mathbf{x})\equiv j\,(\textsf{mod }p)\}-p^{n-1}\big{|}\leq c_{1}p^{n-1}$;
(iii) When $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$,
$\Big{|}\sum\limits_{\mathbf{x}\,(\textsf{mod }p)\atop F(\mathbf{x})\equiv j\,(\textsf{mod }p)}\,\,e_{p}\left(-\mathbf{m}\cdot\mathbf{x}\right)\Big{|}\leq c_{2}p^{\frac{n-1}{2}};$
(iv) When $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$,
#$\#\big{\{}\mathbf{x}(\textsf{mod }p):\,$ $F(\mathbf{x})\equiv j\,(\textsf{mod }p),\,\exists~{}1\leq t\leq n,\,\,s.t.,F_{t}(\mathbf{x})\not\equiv 0\,(\textsf{mod }p),$
$m_{t}\not\equiv 0\,(\textsf{mod }p),m_{i}F_{t}(\mathbf{x})\equiv m_{t}F_{i}(\mathbf{x})\,(\textsf{mod }p)\,(1\leq i\leq n)\big{\}}\leq c_{3}p^{\frac{n-1}{2}}.$
Let $r$ be any positive integer and $E_{1}$, $E_{2}$ be subsets of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$. Assume that
$\delta_{E_{1},E_{2}}\,>\,{\color[rgb]{0,0,0}Cp^{-\frac{n-1}{2}}},$
where $C=(1-c_{1})^{-1}\max\{c_{2},c_{3}\}$. Then
#$\#\{(\mathbf{x},\mathbf{y})\in E_{1}\times E_{2}:\,F(\mathbf{x}-\mathbf{y})\equiv j\,(\textsf{mod }p^{r})\}>0.$
Remark 1.1.
In applications, the constants $c_{1},c_{2},c_{3}$ depend only on the polynomial $F$ (degree, coefficients, number of variables), and the prime $p$ needs to be sufficiently large compared to the degree and the coefficients of $F$. So the result is uniform in all natural numbers $r\geq 1$ and all integers $j$ with $(j,p)=1$. Here and throughout, by depending on $F$ we mean in terms of the degree, coefficients, and number of variables.
Corollary 1.5.
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$, the conclusion of Theorem 1.4 holds for all integers $r,j$ with $r\geq 1$ and $(j,p)=1$, and with the constant $C$ depending only on $F$.
It also worth noting that when $k=3$ and $r=1$, Corollary 1.5 recovers a result established by Iosevich and Koh in [6]. Moreover, Example 7.1 in the last section will show that this distance result is sharp in odd dimensions.
We mention that condition (iii) in Theorem 1.4 is equivalent to
$\Big{|}\sum\limits_{\mathbf{x}\,(\textsf{mod }p)}\sum\limits_{s\not\equiv 0\,(\textsf{mod }p)}\,\,e_{p}\left(-\mathbf{m}\cdot\mathbf{x}+sF(\mathbf{x})-sj\right)\Big{|}\ll p^{\frac{n+1}{2}}$
for $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$. However, the above bound is not applicable to polynomials of the form $F(\mathbf{x})=\sum\nolimits_{j=1}^{n}a_{j}x_{j}^{k_{j}}$ with distinct exponents. In [11], Koh and Shen worked with a weaker exponential sum
$\Big{|}\sum\limits_{\mathbf{x}\,(\textsf{mod }p)}\,\,e_{p}\left(-\mathbf{m}\cdot\mathbf{x}+sF(\mathbf{x})\right)\Big{|}\ll p^{\frac{n}{2}}$
for $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ and $s\not\equiv 0\,(\textsf{mod }p)$. And they proved that $|\Delta_{n,1}(E_{1},E_{2})|\gg p$ whenever $E_{1},E_{2}$ are subsets of $\mathbb{F}_{p}$ satisfying
$\delta_{E_{1},E_{2}}\gg p^{-(n-1)/2}.$ (1)
Merely using the approach in [11] can not lead to a result that is uniform in $r$. With the method developed in this paper, we are able to provide a generalization of the above result in the setting of finite $p$-adic rings.
Theorem 1.6.
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k_{i}}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$, $k_{i}\geq 2$ and $a_{i}\neq 0$ for all $1\leq i\leq n$. Denote $k_{\ast}=\min\nolimits_{1\leq i\leq n}k_{i}$. Then, for any sufficiently large prime $p$, for any natural number $r\geq 1$, and any subsets $E_{1},E_{2}\subseteq\mathbb{Z}/p^{r}\mathbb{Z}$, we have
$|\Delta_{n,r}(E_{1},E_{2})|\gg\min\left\{p^{r},\,\frac{|E_{1}||E_{2}|}{p^{r(2n-2/k_{\ast})-n+1}}\right\}.$
In particular, one has $|\Delta_{n,r}(E_{1},E_{2})|\gg p^{r}$ if
$\delta_{E_{1},E_{2}}\gg p^{r\big{(}\frac{1}{2}-\frac{1}{k_{\ast}}\big{)}-\frac{n-1}{2}}.$ (2)
Here, the implied constants depend only on $F$.
If $k_{\ast}=2$, then we can see that the lower bound of the density is independent of $r$. However, if $k_{\ast}\geq 3$, then the formula (2) is meaningful only when $r\leq\frac{k_{\ast}(n-1)}{k_{\ast}-2}$.
1.2 The conjecture in even dimensions
In even dimensions, one might ask about the conjectured densities to guarantee that $U_{r}\subset\Delta_{n,r}(E)$ or $|\Delta_{n,r}(E)|\gg p^{r}$. We propose the following conjecture.
Conjecture 1.7.
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$ even, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$, if the density $\delta_{E_{1},E_{2}}\gg p^{-\frac{n}{2}}$, then we have $|\Delta_{n,r}(E_{1},E_{2})|\gg p^{r}$.
Example 7.2 in the last section will support this conjecture for the case $k=2$.
In the setting of finite fields, i.e. $r=1$, as mentioned above, Chapman, Erdogan, Hart, Iosevich, and Koh [2] proved the exponent $4/3$ by using an extension theorem associated to circles in the plane. Another proof by using geometric properties of rigid-motions in the plane can be found in [5]. Alex Iosevich asked in several conferences/workshops if the $4/3$-parallel result exists in the setting of finite p-adic rings. In this paper, we give an affirmative answer to his question.
Theorem 1.8.
Let $r\geq 1$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-\frac{2}{3}}$. Then $|\Delta_{2,r}(E)|\gg p^{r}$. Here the implied constant is independent of $r$.
Compared to Theorem 1.4, the exponent $2r-\frac{2}{3}$ gives a smaller threshold density, namely, of $p^{-\frac{2}{3}}$. Regarding the proof of Theorem 1.8, it is challenging if we just want to follow the methods in [2] or in [5]. On the one hand, in a recent exposition, Liao showed that the method in [5] implies too many degenerate cases, even with $r=2$, which are very hard to deal with. On the other hand, a direct computation shows that the same happens with the approach developed in [2], namely, the main difficulty arises when finding the explicit form of the sum $\sum_{j\in\mathbb{Z}/p^{r}\mathbb{Z}}\widehat{1_{C_{r,j}}}(\mathbf{m})\widehat{1_{C_{r,j}}}(\mathbf{m}^{\prime})$ for $\mathbf{m}$ and $\mathbf{m}^{\prime}$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, where $C_{r,j}$ is the circle centered at the origin of radius $j$.
In this paper, to prove Theorem 1.8, we first use a group theoretic argument, which we learned from [19, Appendix], to reduce the theorem to an extension type question for circles, then the rest is devoted to study such type estimates in the p-adic setting. The extension theorems are of independent interest and are expected to have many other applications.
Unlike the finite field case, in finite p-adic rings, we are able to provide a large family of sets $E$ such that Conjecture 1.7 is true, namely, the family of sets $E$ satisfying the property that the density of the fibre of the natural projection over each element in $(\mathbb{Z}/p\mathbb{Z})^{2}$ is not too large in $E$. The precise statement reads as follows.
Theorem 1.9.
Let $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
#$\color[rgb]{0,0,0}\#\{(\mathbf{x}_{1},\mathbf{x}_{2})\in E^{2}:\,\mathbf{x}_{1}\equiv\mathbf{x}_{2}\,(\textsf{mod }p)\}\ll p^{2r-\frac{7}{3}}|E|.$
Then $|\Delta_{2,r}(E)|\gg p^{r}$.
Corollary 1.10.
Let $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
#$\#\{\mathbf{x}^{\prime}\in E\colon\mathbf{x}^{\prime}\equiv\mathbf{x}\,(\textsf{mod }p)\}\ll p^{2r-\frac{7}{3}}$
for each $\mathbf{x}\in E$. Then $|\Delta_{2,r}(E)|\gg p^{r}$.
The implied constants in Theorem 1.9 and Corollary 1.10 are independent of $r$.
1.3 Results on geometric/graph configurations
Let $F(\mathbf{x})$ be a given polynomial in $\mathbb{Z}[\mathbf{x}]$ in $n\geq 2$ variables. We call $F(\mathbf{x})$ good if it satisfies the conditions of Theorem 1.4. In the following, we obtain a number of extensions of the distance result in the setting of geometric/graph configurations in which the polynomials $F(\mathbf{x})$ are assumed to be good.
The first result is on the existence of rectangles of given side-length $j\in U_{r}$, which is an extension of an earlier result due to Lyall and Magyar in [17].
Theorem 1.11.
Let $F$ be a good polynomial. For given $10<\delta^{\prime}<\delta<$ 1 and an integer $j$ with $(j,p)=1$, there exists $\epsilon>0$ such that the following holds. If $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2n}$, with $p^{-\frac{2n-1}{2}}\leq\epsilon$, satisfies $|E|\geq\delta p^{2rn}$, then $E$ contains at least $\delta^{\prime}p^{4rn-2}$ tuples
$(\mathbf{u}_{1},\mathbf{v}_{1}),~{}(\mathbf{u}_{1},\mathbf{v}_{2}),~{}(\mathbf{u}_{2},\mathbf{v}_{1}),~{}(\mathbf{u}_{2},\mathbf{v}_{2})$
that form rectangles of side-length $j$, i.e. $F(\mathbf{u}_{1}-\mathbf{u}_{2})=F(\mathbf{v}_{1}-\mathbf{v}_{2})\equiv j\,(\textsf{mod }p)$.
If we only want to count the number of quadruples $(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4})$ in $E$ that form a cycle, i.e., $F(\mathbf{u}_{1}-\mathbf{u}_{2})=F(\mathbf{u}_{2}-\mathbf{u}_{3})=F(\mathbf{u}_{3}-\mathbf{u}_{4})=F(\mathbf{u}_{4}-\mathbf{u}_{1})\equiv j\,(\textsf{mod }p)$, then a weaker condition is sufficient.
Theorem 1.12.
Let $F$ be a good polynomial. There is some $C>0$ such that, if $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains cycles of length $4$ with distinct vertices and of side-length $j$.
The next result is on the existence of $k$-chains, i.e., a graph of $k+1$ vertices $\mathbf{u}_{1},\ldots,\mathbf{u}_{k+1}$ such that $F(\mathbf{u}_{i}-\mathbf{u}_{i+1})\equiv j\,(\textsf{mod }p)$ for all $11\leq i\leq k-$ 1.
Theorem 1.13.
Let $F$ be a good polynomial, and $k\geq 1$ be an integer. There is some $C>0$ such that, if $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains $k$-chains of side-length $j$.
We note that this theorem improves the density of $p^{-\frac{n-1}{3}}$ due to Lichtin in a recent paper [16].
Our last result is on the distribution of pinned trees. We first need to introduce some notations.
Let $T$ be an arbitrary tree with $k+1$ vertices and $k$ edges. Assume that $V(T)=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{k+1}\}\subset E$, then the edge set of $T$ can be ordered as follows:
$\mathcal{E}(T)=\{(\mathbf{v}_{i_{1}},\mathbf{v}_{i_{2}}),(\mathbf{v}_{i_{3}},\mathbf{v}_{i_{4}}),\ldots,(\mathbf{v}_{i_{2k-1}},\mathbf{v}_{i_{2k}})\},$
where $i_{1}\leq i_{3}\leq\cdots\leq i_{2k-1}$, and $i_{2s}<i_{2t}$ if both $s<t$ and $i_{2s-1}=i_{2t-1}$.
For such a tree $T$, the vector
$\left(|\mathbf{v}_{i_{1}}-\mathbf{v}_{i_{2}}|,|\mathbf{v}_{i_{3}}-\mathbf{v}_{i_{4}}|,\ldots,|\mathbf{v}_{i_{2k-1}}-\mathbf{v}_{i_{2k}}|\right)\in(\mathbb{Z}/p^{r}\mathbb{Z})^{k}$
is called the edge-length vector of $T$. Two trees $T$ and $T^{\prime}$ are called distinct if the corresponding edge-length vectors are not the same. We also recall that two trees $T$ and $T^{\prime}$ are isotropic if there exists a bijective map $\varphi$ from $V(T)$ to $V(T^{\prime})$ such that the edges are preserved. For $\mathbf{x}\in V(T^{\prime})$ and $\mathbf{v}\in V(T)$, we say $(T^{\prime},\mathbf{x})$ is isomorphic to $(T,\mathbf{v})$ if $T$ is isomorphic to $T^{\prime}$ under $\varphi$ and $\varphi(\mathbf{x})=\mathbf{v}$.
Theorem 1.14.
Let $F$ be a good polynomial. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, $k\geq 1$ be an integer, and $(T,\mathbf{v})$ be a given tree with $k+1$ vertices and $k$ edges. If $\delta_{E}\gg_{k}p^{-\frac{n-1}{2}}$, then there exists $\mathbf{x}\in E$ such that the number of distinct trees $(T^{\prime},\mathbf{x})$ with vertices in $E$ and isotropic to $(T,\mathbf{v})$ is $\gg p^{rk}$.
In the rest of the paper, in some places, we may denote $q=p^{r}$ for simplicity if it causes no harm to the argument.
2 Notations and Lemmas
In this paper, the letter $p$ always denote a given prime greater than $2$. The letter $n$ denotes the dimension of the space under consideration, and $r$ is always a positive integer. The letters $m,k,l,i,j$ always denote integers. The imaginary unit is denoted by $\mathsf{i}=\sqrt{-1}$. The cardinality of a finite set $S$ is denoted by either $|S|$ or #$\#S$.
[…]
For any $j\in\mathbb{Z}/p^{r}\mathbb{Z}$, we use $C_{n,r,j}$ to denote the sphere in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ centered at origin with radius $j$, i.e.,
$C_{n,r,j}=\{\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}:\,\|\mathbf{z}\|=j\}.$
Similarly, for a given polynomial $F(x)$ in $n$ variables with coefficients in $\mathbb{Z}$, we define
$S_{n,r,j}=\{\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}:\,F(\mathbf{z})=j\}.$
When it makes no confusion, we abbreviate $C_{n,r,j}$ (or $S_{n,r,j}$) as $C_{r,j}$ (or $S_{r,j}$, respectively).
The additive character modulo $p^{r}$ is denoted by $e_{p^{r}}(x)=e^{\frac{2\pi\mathsf{i}x}{p^{r}}}$, $(x\,\textsf{mod }p^{r})$. For a function $f:\,(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\rightarrow\mathbb{C}$, the Fourier transformation is defined by
$\widehat{f}(\mathbf{m})=\frac{1}{p^{rn}}\sum\limits_{\mathbf{m}\,(\textsf{mod }p^{r})}f(\mathbf{x})e_{p^{r}}(-\mathbf{m}\cdot\mathbf{x}),\quad\big{(}\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\big{)},$
and the Fourier inverse is given by
$f(\mathbf{x})=\sum\limits_{\mathbf{m}\,(\textsf{mod }p^{r})}\widehat{f}(\mathbf{m})e_{p^{r}}(\mathbf{m}\cdot\mathbf{x}),\quad\big{(}\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}).$
The convolution of two functions $f_{1}$ and $f_{2}$ is defined by
$(f_{1}\ast f_{2})(\mathbf{x})=\frac{1}{p^{rn}}\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{r})}f_{1}(\mathbf{x}-\mathbf{y})f_{2}(\mathbf{y}).$
We have the property that $\widehat{f_{1}\ast f_{2}}=\widehat{f_{1}}\widehat{f_{2}}$. The Parseval’s identity is
$\frac{1}{p^{rn}}\sum\limits_{\mathbf{x}\,(\textsf{mod }p^{r})}f_{1}(\mathbf{x})\overline{f_{2}(\mathbf{x})}=\sum\limits_{\mathbf{m}\,(\textsf{mod }p^{r})}\widehat{f_{1}}(\mathbf{m})\overline{\widehat{f_{2}}(\mathbf{m})}.$
The following is one version of Hensel’s lemma that we will regularly applied in this paper.
Lemma 2.1 (Hensel’s lemma).
Let
$\mathbf{G}(\mathbf{x})=\big{(}G_{1}(x_{1},\ldots,x_{n}),\ldots,G_{m}(x_{1},\ldots,x_{n})\big{)}$
be a map from $\mathbb{Z}^{n}$ to $\mathbb{Z}^{m}$, with $G_{i}$ polynomials with integer coefficients. Let $l$ be a positive integer and $\mathbf{y}\in\mathbb{Z}^{n}$. Suppose that $\mathbf{G}(\mathbf{y})\equiv\mathbf{0}$ (mod $(\textsf{mod }p^{l})$. Let $R$ be the rank of $J({\mathbf{G}})|_{\mathbf{y}}$ modulo $p$, where $J({\mathbf{G}})|_{\mathbf{y}}$ is the Jocobi matrix
$J({\mathbf{G}})|_{\mathbf{y}}=\begin{bmatrix}\frac{\partial G_{1}}{\partial x_{1}}(\mathbf{y})&\frac{\partial G_{1}}{\partial x_{2}}(\mathbf{y})&\ldots&\frac{\partial G_{1}}{\partial x_{n}}(\mathbf{y})\\ \frac{\partial G_{2}}{\partial x_{1}}(\mathbf{y})&\frac{\partial G_{2}}{\partial x_{2}}(\mathbf{y})&\ldots&\frac{\partial G_{2}}{\partial x_{n}}(\mathbf{y})\\ \ldots&\ldots&\ldots&\ldots\\ \frac{\partial G_{m}}{\partial x_{1}}(\mathbf{y})&\frac{\partial G_{m}}{\partial x_{2}}(\mathbf{y})&\ldots&\frac{\partial G_{m}}{\partial x_{n}}(\mathbf{y})\end{bmatrix}.$
Then
#$\#\big{\{}\mathbf{z}\,(\textsf{mod }p^{k}):\,\mathbf{G}(\mathbf{y}+p^{l}\mathbf{z})\equiv\mathbf{0}\,(\textsf{mod }p^{l+k})\big{\}}\leq p^{k(n-R)}$ (3)
for any integer $k\geq 1$. When $R=m$, the $\leq$” can be replaced by “$=$”.
Proof.
Firstly, let us consider the case $k=1$. Let $\mathbf{G}(\mathbf{y})=p^{l}\mathbf{b}$, where $\mathbf{b}=(b_{1},b_{2},\ldots,b_{m})\in\mathbb{Z}^{m}$. Note that
$G_{i}(\mathbf{y}+p^{l}\mathbf{z})\equiv G_{i}(\mathbf{y})+p^{l}(\nabla G_{i})(\mathbf{y})\cdot\mathbf{z}\quad(\textsf{mod }p^{l+1})$
for any $i=1,2,\ldots,m$. Then $G_{i}(\mathbf{y}+p^{l}\mathbf{z})\equiv 0\,(\textsf{mod }p^{l+1})$ $(i=1,2,\ldots,m)$ if and only if $b_{i}+(\nabla G_{i})(\mathbf{y})\cdot\mathbf{z}\equiv 0\,(\textsf{mod }p)$ $(i=1,2,\ldots,m)$, if and only if $J_{\mathbf{G}}(\mathbf{y})\mathbf{z}\equiv-\mathbf{b}\,(\textsf{mod }p)$. The number of solutions to such a system of linear equations does not exceed $p^{n-R}$. Thus,
#$\#\big{\{}\mathbf{z}\,(\textsf{mod }p):\,\mathbf{G}(\mathbf{y}+p^{l}\mathbf{z})\equiv\mathbf{0}\,(\textsf{mod }p^{l+1})\big{\}}\leq p^{n-R}.$
Moreover, the system is consistent when $R=m$, and has exactly $p^{n-m}$ solutions.
Secondly, let us consider the case $k=2$. For any $\mathbf{z}_{1}(\textsf{mod }p)$ with $\mathbf{G}(\mathbf{y}+p^{l}\mathbf{z}_{1})\equiv\mathbf{0}\,(\textsf{mod }p^{l+1})$, we write $\mathbf{y}_{1}=\mathbf{y}+p^{l}\mathbf{z}_{1}$. Noting that $J(\mathbf{G})|_{\mathbf{y}_{1}}\equiv J(\mathbf{G})|_{\mathbf{y}}\,(\textsf{mod }p)$, the rank of $J(\mathbf{G})|_{\mathbf{y}_{1}}$ modulo $p$ is also $R$. Hence
#$\#\big{\{}\mathbf{z}_{2}\,(\textsf{mod }p):\,\mathbf{G}(\mathbf{y}_{1}+p^{l+1}\mathbf{z}_{2})\equiv\mathbf{0}\,(\textsf{mod }p^{l+2})\big{\}}\leq p^{n-R}.$
It follows that
#$\#\big{\{}(\mathbf{z}_{1},\mathbf{z}_{2})\,(\textsf{mod }p):\,\mathbf{G}\big{(}\mathbf{y}+p^{l}(\mathbf{z}_{1}+p\mathbf{z}_{2})\big{)}\equiv\mathbf{0}\,(\textsf{mod }p^{l+2})\big{\}}\leq p^{2(n-R)}.$
Finally, the conclusion follows by an induction on $k$. ∎
Lemma 2.2 (Weil’s theorem).
Let $f(x)\in\mathbb{F}_{p}[x]$ be a polynomial of degree $k\geq 1$ with $(k,p)=1$. Then
$\Big{|}\sum\limits_{x\in\mathbb{F}_{p}}e_{p}(f(x))\Big{|}\leq(k-1)p^{1/2}.$
Lemma 2.3 (Theorem 2.3, [11]).
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k_{i}}$ be a polynomial in $\mathbb{F}_{p}[\mathbf{x}]$ such that $n\geq 2$, $k_{i}\geq 1$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Let $j\in\mathbb{F}_{p}^{\ast}$. Then
#$\#\{\mathbf{x}\in\mathbb{F}_{p}:\,F(\mathbf{x})=j\}=(1+\textit{o}_{F}(1))p^{n-1}$
as $p\rightarrow\infty$.
Lemma 2.4 (Lemma 2.1, [11]).
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{F}_{p}[\mathbf{x}]$ with $n\geq 1$, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Let $j\in\mathbb{F}_{p}^{\ast}$. Then, for any $\mathbf{m}\in\mathbb{F}_{p}^{n}\setminus\{\mathbf{0}\}$, one has
$\sum\limits_{\mathbf{x}\in\mathbb{F}_{p}^{n}\atop F(\mathbf{x})=j}e_{p}(-\mathbf{m}\cdot\mathbf{x})\ll_{n,k}p^{\frac{n-1}{2}}$
when $p$ is sufficiently large.
Lemma 2.5.
Let $p$ be a prime and $j$ be an integer with $(j,p)=1$. Assume that $F(\mathbf{x})$ is a polynomial in $\mathbb{Z}[\mathbf{x}]$ in $n\geq 1$ variables satisfying the conditions (i) and (iii) in Theorem 1.4, and the following:
(iv’): When $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$,
#$\#\big{\{}\mathbf{y}(\textsf{mod }p):\,F(\mathbf{y})\equiv j(\textsf{mod }p),\,\exists~{}1\leq t\leq n,\,\,s.t.,F_{t}(\mathbf{y})\not\equiv 0(\textsf{mod }p),$
$\qquad\qquad m_{t}\not\equiv 0(\textsf{mod }p),m_{i}F_{t}(\mathbf{y})\equiv m_{t}F_{i}(\mathbf{y})\,(\textsf{mod }p)\,\mbox{for all}~{}1\leq i\leq n\big{\}}\leq{c_{3}}p^{\kappa},$
for some $c_{3}>0$ and $0\leq\kappa\leq\frac{n-1}{2}$.
Then, for any $r\geq 1$, we have
$|\widehat{1_{S_{r,j}}}(\mathbf{m}^{\prime})|\leq\begin{cases}c_{2}\,p^{-r-\frac{n-1}{2}},\quad&\text{if }\mathbf{m}^{\prime}\equiv\mathbf{0}\,(\textsf{mod }p^{r-1})\text{ but }\mathbf{m}^{\prime}\not\equiv\mathbf{0}\,(\textsf{mod}p^{r}),\\ c_{3}\,p^{-r-n+1+\kappa},&\text{if }\mathbf{m}^{\prime}\not\equiv\mathbf{0}\,(\textsf{mod }p^{r-1}).\end{cases}$
Proof.
Recall that $S_{r,j}=\{\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\colon F(\mathbf{x})\equiv j\,(\textsf{mod }p^{r})\}$. For simplicity, we denote $v_{\mathbf{m}^{\prime}}=\nu$. Then $\mathbf{m}^{\prime}=p^{\nu}\mathbf{m}$ with $\mathbf{m}\in(\mathbb{Z}/p^{r-\nu}\mathbb{Z})^{n}$ and $v_{\mathbf{m}}=0$.
When $r=1$, the conclusion follows from the condition (iii). In the following, we assume that $r\geq 2$.
For $11\leq\nu\leq r-$ 1, by a change of variables $\mathbf{x}=\mathbf{y}+p^{r-\nu}\mathbf{z}$, we have that
$\widehat{1_{S_{r,j}}}(\mathbf{m}^{\prime})=$ $\frac{1}{p^{rn}}\sum\limits_{\mathbf{x}(\textsf{mod }p^{r})\atop F(\mathbf{x})\equiv j(\textsf{mod }p^{r})}e_{p^{r}}(-\mathbf{m}^{\prime}\cdot\mathbf{x})$
$=\frac{1}{p^{rn}}\sum\limits_{\mathbf{y}(\textsf{mod }p^{r-\nu})\atop F(\mathbf{y})\equiv j(\textsf{mod }p^{r-\nu})}\sum\limits_{\mathbf{z}(\textsf{mod }p^{\nu})\atop F(\mathbf{y}+p^{r-\nu}\mathbf{z})\equiv j(\textsf{mod }p^{r})}e_{p^{r}}\big{(}-(p^{\nu}\mathbf{m})\cdot(\mathbf{y}+p^{r-\nu}\mathbf{z})\big{)}$
$=\frac{1}{p^{rn}}\sum\limits_{\mathbf{y}(\textsf{mod }p^{r-\nu})\atop F(\mathbf{y})\equiv j(\textsf{mod }p^{r-\nu})}e_{p^{r-\nu}}\big{(}-\mathbf{m}\cdot\mathbf{y})\sum\limits_{\mathbf{z}(\textsf{mod }p^{\nu})\atop F(\mathbf{y}+p^{r-\nu}\mathbf{z})\equiv j(\textsf{mod }p^{r})}1.$
Since $F(\mathbf{y})\equiv j\,(\textsf{mod }p)$, one has $(\nabla F)(\mathbf{y})\not\equiv\mathbf{0}\,(\textsf{mod }p)$ by the condition (i). By Hensel’s lemma, we have
$\sum\limits_{\mathbf{z}(\textsf{mod }p^{\nu})\atop F(\mathbf{y}+p^{r-\nu}\mathbf{z})\equiv j(\textsf{mod }p^{r})}1=p^{\nu(n-1)}$
[…]
$\widehat{1_{S_{r,j}}}(\mathbf{m}^{\prime})=p^{-rn+\nu(n-1)}\sum\limits_{\mathbf{y}(\textsf{mod }p^{r-\nu})\atop F(\mathbf{y})\equiv j(\textsf{mod }p^{r-\nu})}e_{p^{r-\nu}}\big{(}-\mathbf{m}\cdot\mathbf{y})$ (4)
for $11\leq\nu\leq r-$ 1. Moreover, when $\nu=0$, the expression (4) trivially holds.
Let us substitute the parameters $\mathbf{x}$, $\gamma$ for $\mathbf{y}$, $r-\nu$, respectively. One has $1\leq\gamma\leq r$. Then, it is sufficient to show that
$T_{\gamma}(\mathbf{m}):=\sum\limits_{\mathbf{x}(\textsf{mod }p^{\gamma})\atop F(\mathbf{x})\equiv j(\textsf{mod }p^{\gamma})}e_{p^{\gamma}}\big{(}-\mathbf{m}\cdot\mathbf{x})\leq\begin{cases}c_{2}\,p^{(\gamma-1/2)(n-1)},\quad&\text{if }{\color[rgb]{0,0,0}\gamma=1},\\ c_{3}\,p^{(\gamma-1)(n-1)+\kappa},&\text{if }{\color[rgb]{0,0,0}2\leq\gamma\leq r}.\end{cases}$
Here $v_{\mathbf{m}}=0$, i.e., $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$.
When $\gamma=1$, the bound follows from the condition (iii). In the following, we consider the situation that $\gamma\geq 2$. By a change of variables $\mathbf{x}=\mathbf{y}+p^{\gamma-1}\mathbf{z}$,
$T_{\gamma}(\mathbf{m})$ $=\sum\limits_{\mathbf{y}(\textsf{mod }p^{\gamma-1})\atop F(\mathbf{y})\equiv j(\textsf{mod }p^{\gamma-1})}\sum\limits_{\mathbf{z}(\textsf{mod }p)\atop F(\mathbf{y}+p^{\gamma-1}\mathbf{z})\equiv j(\textsf{mod }p^{\gamma})}e_{p^{\gamma}}\big{(}-\mathbf{m}\cdot(\mathbf{y}+p^{\gamma-1}\mathbf{z}))$
$=\sum\limits_{\mathbf{y}(\textsf{mod }p^{\gamma-1})\atop F(\mathbf{y})\equiv j(\textsf{mod }p^{\gamma-1})}e_{p^{\gamma}}\big{(}-\mathbf{m}\cdot\mathbf{y})\sum\limits_{\mathbf{z}(\textsf{mod }p)\atop F(\mathbf{y}+p^{\gamma-1}\mathbf{z})\equiv j(\textsf{mod }p^{\gamma})}e_{p}\big{(}-\mathbf{m}\cdot\mathbf{z}).$
Let us denote
$T_{\gamma}(\mathbf{m};\mathbf{y}):=\sum\limits_{\mathbf{z}(\textsf{mod }p)\atop F(\mathbf{y}+p^{\gamma-1}\mathbf{z})\equiv j(\textsf{mod }p^{\gamma})}e_{p}\big{(}-\mathbf{m}\cdot\mathbf{z}).$
The condition $F(\mathbf{y}+p^{\gamma-1}\mathbf{z})\equiv j(\textsf{mod }p^{\gamma})$ is equivalent to
$F(\mathbf{y})+p^{\gamma-1}(\nabla F)(\mathbf{y})\cdot\mathbf{z}\equiv j\,(\textsf{mod }p^{\gamma}).$
Suppose that $F(\mathbf{y})\equiv j+p^{\gamma-1}h_{\gamma}(\mathbf{y})$ (mod $(\textsf{mod }p^{\gamma})$ for some $h_{\gamma}(\mathbf{y})$ (mod $(\textsf{mod }p)$. Since $\gamma\geq 2$ and $F(\mathbf{y})\equiv j\,(\textsf{mod }p^{\gamma-1})$, one has $(\nabla F)(\mathbf{y})\not\equiv\mathbf{0}\,(\textsf{mod }p)$ by the condition (i). Recall that $F_{i}(\mathbf{y})=\frac{\partial F}{\partial x_{i}}(\mathbf{y})$. Without loss of generality, let us proceed with $F_{1}(\mathbf{y})\not\equiv 0(\textsf{mod }p)$. Then
$z_{1}\equiv-F_{1}^{-1}(\mathbf{y})(h_{\gamma}(\mathbf{y})+F_{2}(\mathbf{y})z_{2}+\ldots+F_{n}(\mathbf{y})z_{n})\quad(\textsf{mod }p).$
Here $F_{1}^{-1}(\mathbf{y})$ is the inverse of $F_{1}(\mathbf{y})$ (mod $(\textsf{mod }p)$. So
$T_{\gamma}(\mathbf{m};\mathbf{y})=e_{p}\big{(}m_{1}F_{1}^{-1}(\mathbf{y})h_{\gamma}(\mathbf{y}))\prod_{i=2}^{n}\sum\limits_{z_{i}(\textsf{mod }p)}e_{p}\Big{(}\big{(}-m_{i}+m_{1}F_{1}^{-1}(\mathbf{y})F_{i}(\mathbf{y})\big{)}z_{i}\Big{)}.$
It follows that
$|T_{\gamma}(\mathbf{m};\mathbf{y})|=\begin{cases}p^{n-1},\quad&\text{if }m_{i}F_{1}(\mathbf{y})\equiv m_{1}F_{i}(\mathbf{y})\,\,(\textsf{mod }p)\text{ for }1\leq i\leq n,\\ 0,\quad&\text{otherwise}.\end{cases}$
(Note that, for $i=1$, the formula $m_{i}F_{1}(\mathbf{y})\equiv m_{1}F_{i}(\mathbf{y})\,\,(\textsf{mod }p)$ trivially holds.)
Moreover, if $m_{1}\equiv 0$ (mod $(\textsf{mod }p)$, then the assumption $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ shows that $m_{i_{0}}\not\equiv 0\,(\textsf{mod }p)$ for some $2\leq i_{0}\leq n$. Then $m_{i_{0}}F_{1}(\mathbf{y})\not\equiv m_{1}F_{i_{0}}(\mathbf{y})\,(\textsf{mod }p)$, which leads to $T_{\gamma}(\mathbf{m};\mathbf{y})=0$. As a result, if $T(\mathbf{m};\mathbf{y})\neq 0$ and $F_{t}(\mathbf{y})\not\equiv 0$ (mod $(\textsf{mod }p)$ for some $1\leq t\leq n$, then $m_{t}\not\equiv 0$ (mod $(\textsf{mod }p)$.
Now we conclude that
$|T_{\gamma}(\mathbf{m})|\leq p^{n-1}|V_{\gamma-1}(\mathbf{m})|,$
where
$V_{\mu}(\mathbf{m}):=$$\big{\{}\mathbf{y}(\textsf{mod }p^{\mu}):\,F(\mathbf{y})\equiv j(\textsf{mod }p^{\mu}),\,\exists~{}1\leq t\leq n,\,\,s.t.,F_{t}(\mathbf{y})\not\equiv 0(\textsf{mod }p),$
$\qquad\qquad m_{t}\not\equiv 0(\textsf{mod }p),m_{i}F_{t}(\mathbf{y})\equiv m_{t}F_{i}(\mathbf{y})(\textsf{mod }p)\,\mbox{for all}~{}1\leq i\leq n\big{\}}.$
When $\gamma\geq 3$, we further have
$|V_{\gamma-1}(\mathbf{m})|\leq$ $\sum\limits_{\mathbf{w}\in V_{1}(\mathbf{m})}\#\{\mathbf{y}(\textsf{mod }p^{\gamma-1}):\,F(\mathbf{y})\equiv j(\textsf{mod }p^{\gamma-1}),\,\mathbf{y}\equiv\mathbf{w}\,(\textsf{mod }p)\}.$
Note that $F(\mathbf{w})\equiv j\,(\textsf{mod }p)$ for any $\mathbf{w}\in V_{1}(\mathbf{m})$. So $(\nabla F)(\mathbf{w})\not\equiv 0\,(\textsf{mod }p)$ by the condition (i). Applying Hensel’s lemma again, we have
#$\#\{\mathbf{y}(\textsf{mod }p^{\gamma-1}):\,F(\mathbf{y})\equiv j(\textsf{mod }p^{\gamma-1}),\,\mathbf{y}\equiv\mathbf{w}\,(\textsf{mod }p)\}$
=#$=\#\{\mathbf{z}\,(\textsf{mod }p^{\gamma-2}):\,F(\mathbf{w}+p\mathbf{z})\equiv j\,(\textsf{mod }p^{\gamma-1})\}=p^{(\gamma-2)(n-1)}$
for each given $\mathbf{w}$. So $|V_{r-1}(\mathbf{m})|\leq p^{(\gamma-2)(n-1)}|V_{1}(\mathbf{m})|$. It follows that, for any $\gamma\geq 2$,
$|T_{\gamma}(\mathbf{m})|\leq p^{(\gamma-1)(n-1)}|V_{1}(\mathbf{m})|.$
Recalling that $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$, one can apply the condition (iv’) to get $|V_{1}(\mathbf{m})|\leq c_{3}p^{\kappa}$. Thus, we conclude that $|T_{\gamma}(\mathbf{m})|\leq c_{3}p^{(\gamma-1)(n-1)+\kappa}$. The proof is completed. ∎
Remark 2.1.
[…]
Remark 2.2.
When $n=1$, the cardinality of the set in condition (iv’) becomes
#$\#\{y\,(\textsf{mod }p):\,F(y)\equiv j\,(\textsf{mod }p)\},$
which is no larger than the degree of $F(y)$. So one may take $\kappa=0$ in this case.
Lemma 2.6.
Let $\kappa$ be a number with $20\leq\kappa\leq n/$ 2 and $c_{1},c_{2}>0$ be constants. Let $F(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ be a polynomial in $n\geq 1$ variables of degree $\geq 2$ such that the upper bounds
$\sum\limits_{\mathbf{x}\,(\textsf{mod }p)}e_{p}\left(sF(\mathbf{x})+\mathbf{m}\cdot\mathbf{x}\right)\leq c_{1}p^{n/2}$ (5)
and
#$\#\{\mathbf{x}\,(\textsf{mod }p):\,s\,(\nabla F)(\mathbf{x})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)\}\leq c_{2}p^{\kappa}$ (6)
both hold when $(\mathbf{m},s)\not\equiv(\mathbf{0},0)\,(\textsf{mod }p)$. Then, for any $r\geq 1$, we have
$\sum\limits_{\mathbf{x}\,(\textsf{mod }p^{r})}e_{p^{r}}\left(s^{\prime}F(\mathbf{x})+\mathbf{m}^{\prime}\cdot\mathbf{x}\right)$
$\leq\begin{cases}c_{1}p^{(r-1/2)n},\quad&\text{if }(\mathbf{m}^{\prime},\,s^{\prime})\equiv(\mathbf{0},0)\,(\textsf{mod }p^{r-1})\text{ but }(\mathbf{m}^{\prime},\,s^{\prime})\not\equiv(\mathbf{0},0)\,(\textsf{mod }p^{r}),\\ c_{2}p^{(r-1)n+\kappa},&\text{if }(\mathbf{m}^{\prime},\,s^{\prime})\not\equiv(\mathbf{0},0)\,(\textsf{mod }p^{r-1}).\end{cases}$
Proof.
For simplicity, let us write $v_{(\mathbf{m}^{\prime},s^{\prime})}=\nu$. Since $(\mathbf{m}^{\prime},\,s^{\prime})\not\equiv(\mathbf{0},0)\,(\textsf{mod }p^{r})$, we have $10\leq\nu\leq r-$ 1. Write $\mathbf{m}^{\prime}=p^{\nu}\mathbf{m}$ and $s^{\prime}=p^{\nu}s$. Then $(\mathbf{m},s)\not\equiv(\mathbf{0},0)\,(\textsf{mod }p)$.
When $11\leq\nu\leq r-$ 1, we obtain by a change of variables $\mathbf{x}=\mathbf{y}+p^{r-\nu}\mathbf{z}$ that
$G_{r}(\mathbf{m}^{\prime},s^{\prime}):=$ $\sum\limits_{\mathbf{x}\,(\textsf{mod }p^{r})}e_{p^{r}}\left(s^{\prime}F(\mathbf{x})+\mathbf{m}^{\prime}\cdot\mathbf{x}\right)$
$=\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{r-\nu})}\sum\limits_{\mathbf{z}\,(\textsf{mod }p^{\nu})}e_{p^{r}}\left(p^{\nu}sF(\mathbf{y})+p^{\nu}s(\nabla F)(\mathbf{y})\cdot(p^{r-\nu}\mathbf{z})+p^{\nu}\mathbf{m}\cdot(\mathbf{y}+p^{r-\nu}\mathbf{z})\right)$
$=\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{r-\nu})}e_{p^{r-\nu}}\left(sF(\mathbf{y})+\mathbf{m}\cdot\mathbf{y}\right)\sum\limits_{\mathbf{z}\,(\textsf{mod }p^{\nu})}1=p^{\nu n}\cdot G_{r-\nu}(\mathbf{m},s).$
When $\nu=0$, the above equality trivially holds. By writing $\gamma=r-\nu$ with $1\leq\gamma\leq r$, it is sufficient to prove that
$G_{\gamma}(\mathbf{m},s)\leq\begin{cases}c_{1}p^{(\gamma-1/2)n},\quad&\text{if}{\color[rgb]{0,0,0}\gamma=1},\\ c_{2}p^{(\gamma-1)n+\kappa},&\text{if }{\color[rgb]{0,0,0}2\leq\gamma\leq r},\end{cases}$
where $(\mathbf{m},s)\not\equiv(\mathbf{0},0)\,(\textsf{mod }p)$.
When $\gamma=1$, the conclusion follows from (5). In the following, let us assume that $\gamma\geq 2$. By a change of variables $\mathbf{x}=\mathbf{y}+p^{\gamma-1}\mathbf{z}$, we obtain that
$G_{\gamma}(\mathbf{m},s)$ $=\sum\limits_{\mathbf{x}\,(\textsf{mod }p^{\gamma})}e_{p^{\gamma}}\left(sF(\mathbf{x})+\mathbf{m}\cdot\mathbf{x}\right)$
$=\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{\gamma-1})}e_{p^{\gamma}}\left(sF(\mathbf{y})+\mathbf{m}\cdot\mathbf{y}\right)\sum\limits_{\mathbf{z}\,(\textsf{mod }p)}e_{p}\big{(}(s\,(\nabla F)(\mathbf{y})+\mathbf{m})\cdot\mathbf{z}\big{)}$
$=p^{n}\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{\gamma-1})\atop s(\nabla F)(\mathbf{y})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)}e_{p^{\gamma}}\left(sF(\mathbf{y})+\mathbf{m}\cdot\mathbf{y}\right)$
$\leq p^{n}\cdot\#\{\mathbf{y}\,(\textsf{mod }p^{\gamma-1}):\,s\,(\nabla F)(\mathbf{y})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)\}.$
When $\gamma=2$, it follows from (6) that $G_{2}(\mathbf{m},s)\leq c_{2}p^{n+\kappa}$. When $\gamma\geq 3$, we further have
#$\#\{\mathbf{y}\,(\textsf{mod }p^{\gamma-1}):\,s\,(\nabla F)(\mathbf{y})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)\}$
=#$=\#\{(\mathbf{w},\mathbf{z})\in(\mathbb{Z}/p\mathbb{Z})^{n}\times(\mathbb{Z}/p^{\gamma-2}\mathbb{Z})^{n}:\,s\,(\nabla F)(\mathbf{w}+p\mathbf{z})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)\}$
$=p^{(\gamma-2)n}\cdot\#\{\mathbf{w}\,(\textsf{mod }p):\,s\,(\nabla F)(\mathbf{w})+\mathbf{m}\equiv\mathbf{0}\,(\textsf{mod }p)\}.$
By (6) again, one deduces that
$G_{\gamma}(\mathbf{m},s)\leq p^{n}\cdot p^{(\gamma-2)n}\cdot c_{2}p^{\kappa}=c_{2}p^{(\gamma-1)n+\kappa}.$
The proof is completed. ∎
Lemma 2.7.
Let $a$ be an integer with $(a,p)=1$. Let $k\geq 2$ be a natural number with $(k,p)=1$. Then for any $m\,(\textsf{mod }p)$,
$\sum\limits_{s\,(\textsf{mod }p)}\left|\sum\limits_{x\,(\textsf{mod }p)}e_{p}(-mx+sax^{k})\right|^{2}\ll_{k}p^{2}.$ (7)
Moreover, for any natural number $r\geq 2$, it satiesfies that
$\sum\limits_{s\,(\textsf{mod }p^{r})}\left|\sum\limits_{x\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx+sax^{k})\right|^{2}\ll_{k}\begin{cases}p^{2r},\quad&\text{if }{\color[rgb]{0,0,0}\textsf{ord}_{p}}(m)<r-\lceil\frac{r}{k}\rceil,\\ p^{(3-2/k)r},&\text{if }{\color[rgb]{0,0,0}\textsf{ord}_{p}}(m)\geq r-\lceil\frac{r}{k}\rceil,\end{cases}$ (8)
for $m\in\mathbb{Z}/p^{r}\mathbb{Z}$.
Proof.
Note that
$\sum\limits_{s\,(\textsf{mod }p^{r})}\left|\sum\limits_{x\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx+sax^{k})\right|^{2}=\sum\limits_{x,y\,(\textsf{mod }p^{r})}\sum\limits_{s\,(\textsf{mod }p^{r})}e_{p^{r}}\big{(}-m(x-y)+sa(x^{k}-y^{k})\big{)}$
$\qquad=p^{r}\sum\limits_{x,y\,(\textsf{mod }p^{r})\atop x^{k}\equiv y^{k}\,(\textsf{mod }p^{r})}e_{p^{r}}\big{(}-m(x-y)\big{)}=p^{r}\sum\limits_{j\,(\textsf{mod }p^{r})}\Big{|}\sum\limits_{x\,(\textsf{mod }p^{r})\atop x^{k}\equiv j\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx)\Big{|}^{2}.$
When $j\not\equiv 0\,(\textsf{mod }p)$, one has
$\Big{|}\sum\limits_{x\,(\textsf{mod }p^{r})\atop x^{k}\equiv j\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx)\Big{|}\leq\#\{x\,(\textsf{mod }p^{r}):\,x^{k}\equiv j\,(\textsf{mod }p^{r})\}.$ (9)
When $r=1$, it is not hard to see that the above quantity is $\ll_{k}1$. By Hensel’s lemma, we conclude that (9) is $\ll_{k}1$ for any $r\geq 1$. Hence
$p^{r}\sum\limits_{j\not\equiv 0\,(\textsf{mod }p^{r})}\Big{|}\sum\limits_{x\,(\textsf{mod }p^{r})\atop x^{k}\equiv j\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx)\Big{|}^{2}\ll p^{2r}.$
Next, we consider the case that $j\equiv 0\,(\textsf{mod }p)$. When $r=1$, it is easy to see that
$p\cdot\Big{|}\sum\limits_{x\,(\textsf{mod }p)\atop x^{k}\equiv 0\,(\textsf{mod}p)}e_{p}(-mx)\Big{|}^{2}=p.$
When $r\geq 2$, let us take $\mu=\lceil\frac{r}{k}\rceil$. Then $x^{k}\equiv 0\,(\textsf{mod }p^{r})$ if and only if $p^{\mu}|x$. By writing $x=p^{\mu}y$, we obtain that
$\sum\limits_{x\,(\textsf{mod }p^{r})\atop x^{k}\equiv 0\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx)=\sum\limits_{y\,(\textsf{mod }p^{r-\mu})}e_{p^{r-\mu}}(-my)\ll p^{r-\mu}\ll p^{r(1-1/k)}.$
Donote $\nu=\textsf{ord}_{p}(m)$ and write $m=p^{\nu}\widetilde{m}$. When $1\leq\nu<r-\mu$, one deduces by a change of variables $y=w+p^{r-\mu-\nu}z$ that
$\sum\limits_{y\,(\textsf{mod }p^{r-\mu})}e_{p^{r-\mu}}(-my)=\sum\limits_{w\,(\textsf{mod }p^{r-\mu-\nu})}e_{p^{r-\mu-\nu}}(-\tilde{m}w)\sum\limits_{z\,(\textsf{mod }p^{\nu})}1=0.$
When $\nu=0$, the above result also holds. So the summand over $j\equiv 0\,(\textsf{mod }p^{r})$ is
$p^{r}\cdot\Big{|}\sum\limits_{x\,(\textsf{mod }p^{r})\atop x^{k}\equiv 0\,(\textsf{mod }p^{r})}e_{p^{r}}(-mx)\Big{|}^{2}\ll\begin{cases}0,\quad&\text{if }\nu<r-\lceil\frac{r}{k}\rceil,\\ p^{r(3-2/k)},&\text{if }\nu\geq r-\lceil\frac{r}{k}\rceil.\end{cases}$
The lemma then follows. ∎
3 Proof of Theorems 1.4, 1.6, and Corollary 1.5
For simplicity, we write $q=p^{r}$ in this section.
Proof of Therorem 1.4.
Recall that $S_{r,j}=\{\mathbf{z}\in(\mathbb{Z}/q\mathbb{Z})^{n}\colon F(\mathbf{z})\equiv j\,(\textsf{mod }q)\}$. First, we have
$\mathcal{N}_{r,j}:=$$\#\{(\mathbf{x},\mathbf{y})\in E_{1}\times E_{2}:\,F(\mathbf{x}-\mathbf{y})\equiv j\,(\textsf{mod }q)\}$
$=\sum\limits_{\mathbf{x},\mathbf{y}\,(\textsf{mod }q)}1_{E_{1}}(\mathbf{x})1_{E_{2}}(\mathbf{y})1_{S_{r,j}}(\mathbf{x}-\mathbf{y})=q^{2n}\cdot\frac{1}{q^{n}}\sum\limits_{\mathbf{x}\,(\textsf{mod }q)}1_{E_{1}}(\mathbf{x})(1_{E_{2}}\ast 1_{S_{r,j}})(\mathbf{x})$
$=q^{2n}\sum\limits_{\mathbf{m}\,(\textsf{mod }q)}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\widehat{1_{E_{1}}\ast 1_{S_{r,j}}}(\mathbf{m})=q^{2n}\sum\limits_{\mathbf{m}\,(\textsf{mod }q)}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\widehat{1_{E_{1}}}(\mathbf{m})\widehat{1_{S_{r,j}}}(\mathbf{m})=\mathcal{M}+\mathcal{E},$
where
$\mathcal{M}=q^{-n}|E_{1}||E_{2}||{S_{r,j}}|~{}~{}\mbox{and}~{}~{}\mathcal{E}=q^{2n}\sum\limits_{\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\widehat{1_{E_{1}}}(\mathbf{m})\widehat{1_{S_{r,j}}}(\mathbf{m}).$
By the condition (ii), one has
|#$|\#S_{1,j}-p^{n-1}|\leq c_{1}p^{n-1}.$
With the condition (i), we may apply Hensel’s lemma to get
|#$\big{|}\#S_{r,j}-q^{n-1}\big{|}\leq c_{1}q^{n-1}$
for all $r,j$ with $r\geq 1$ and $(j,p)=1$.
Applying Lemma 2.5, one obtains that
$\left|\widehat{1_{S_{r,j}}}(\mathbf{m})\right|\leq\max\{c_{2},c_{3}\}\cdot q^{-1}p^{-\frac{n-1}{2}}$
for $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)$. Hence,
$|\mathcal{E}|$ $\leq q^{2n}\cdot\sup\limits_{\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)}\big{|}\widehat{1_{S_{r,j}}}(\mathbf{m})\big{|}\cdot\left(\sum\limits_{\mathbf{m}\,(\textsf{mod }q)}\big{|}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\big{|}^{2}\right)^{1/2}\left(\sum\limits_{\mathbf{m}\,(\textsf{mod }q)}\big{|}\overline{\widehat{1_{E_{2}}}(\mathbf{m})}\big{|}^{2}\right)^{1/2}$
$=q^{n}|E_{1}|^{1/2}|E_{2}|^{1/2}\cdot\sup\limits_{\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)}\big{|}\widehat{1_{S_{r,j}}}(\mathbf{m})\big{|}\leq\max\{c_{2},c_{3}\}\cdot\sqrt{|E_{1}||E_{2}|}\cdot q^{n-1}p^{-(n-1)/2}.$
Combining all bounds, we conclude that
$\left|\mathcal{N}_{r,j}-\frac{|E_{1}||E_{2}|}{q}\right|\leq\frac{c_{1}|E_{1}||E_{2}|}{q}+\max\{c_{2},c_{3}\}\cdot\sqrt{|E_{1}||E_{2}|}\cdot q^{n-1}p^{-(n-1)/2}.$
So $\mathcal{N}_{r,j}>0$ if
$\frac{\sqrt{|E_{1}||E_{2}|}}{q^{n}}>\frac{\max\{c_{2},c_{3}\}}{(1-c_{1})p^{\frac{n-1}{2}}}.$
The proof is completed. ∎
Proof of Corollary 1.5.
Let $p>\max\{k(k-1),|a_{1}|,\ldots,|a_{n}|\}$ be sufficiently large. Then $(k,p)=1$ and $(a_{i},p)=1$ for all $1\leq i\leq n$.
Let us verify the four conditions in Theorem 1.4 for $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ for all $j\in\mathbb{Z}$ with $(j,p)=1$. We have
$(\nabla F)(\mathbf{x})=(ka_{1}x_{1}^{k-1},\ldots,ka_{n}x_{n}^{k-1})\not\equiv\mathbf{0}\,(\textsf{mod }p)$
when $\mathbf{x}\not\equiv\mathbf{0}\,(\textsf{mod }p)$. Then the condition (i) holds. Moreover, the conditions (ii) and (iii) follow from Lemmas 2.3 and 2.4, respectively. In the following, let us verify condition (iv).
For $\mathbf{m}{\color[rgb]{0,0,0}=(m_{1},m_{2},\ldots,m_{n})}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ and $(j,p)=1$, the cardinality of the set in the condition (iv), denoted by $\mathfrak{N}$, satisfies that
$\mathfrak{N}$ $\leq\sum\limits_{1\leq t\leq n\atop m_{t}\not\equiv 0\,(\textsf{mod }p)}\#\left\{\mathbf{x}\,(\textsf{mod }p):\,\sum\limits_{i=1}^{n}a_{i}x_{i}^{k}\equiv j\,(\textsf{mod }p),\right.$
$\qquad\qquad\left.x_{t}\not\equiv 0\,(\textsf{mod }p),\,m_{i}a_{t}x_{t}^{k-1}\equiv m_{t}a_{i}x_{i}^{k-1}\,(\textsf{mod }p)\,\mbox{for all}~{}1\leq i\leq n\right\}.$
Thus, for the $\mathbf{x}$ taken into account, we have $x_{i}^{k-1}\equiv(m_{t}a_{i})^{-1}m_{i}a_{t}x_{t}^{k-1}\,(\textsf{mod }p)$.
When $k=2$, it follows that
$\mathfrak{N}\leq\sum\limits_{1\leq t\leq n\atop m_{t}\not\equiv 0\,(\textsf{mod }p)}\#\left\{x_{t}\,(\textsf{mod }p):\,\Big{(}\sum\limits_{i=1}^{n}\big{(}(m_{t}a_{i})^{-1}m_{i}a_{t}\big{)}^{2}a_{i}\Big{)}x_{t}^{2}\equiv j\,(\textsf{mod }p)\right\}\leq 2n\ll 1.$
When $k=3$, we further have
$x_{i}^{3}\equiv x_{i}^{3+p}\equiv\big{(}(m_{t}a_{i})^{-1}m_{i}a_{t}x_{t}^{2}\big{)}^{\frac{3+p}{2}}\equiv\big{(}(m_{t}a_{i})^{-1}m_{i}a_{t}\big{)}^{\frac{3+p}{2}}x_{t}^{3}\,(\textsf{mod }p).$
So
$\mathfrak{N}\leq\sum\limits_{1\leq t\leq n\atop m_{t}\not\equiv 0\,(\textsf{mod }p)}\#\left\{x_{t}\,(\textsf{mod }p):\,\Big{(}\sum\limits_{i=1}^{n}\big{(}(m_{t}a_{i})^{-1}m_{i}a_{t}\big{)}^{\frac{3+p}{2}}a_{i}\Big{)}x_{t}^{3}\equiv j\,(\textsf{mod }p)\right\}\leq 3n\ll 1.$
When $k\geq 4$, we have
$\mathfrak{N}\leq\sum\limits_{1\leq t\leq n\atop m_{t}\not\equiv 0\,(\textsf{mod }p)}\#\left\{\mathbf{x}\,(\textsf{mod }p):\,x_{i}^{k-1}\equiv(m_{t}a_{i})^{-1}m_{i}a_{t}x_{t}^{k-1}\,(\textsf{mod }p)\,\mbox{for all}~{}1\leq i\leq n\right\}.$
The variable $x_{t}$ can be chosen mod $p$, and each $x_{i}$ $(i\neq t)$ has at most $k-1$ choices. So
$\mathfrak{N}\leq n\cdot p\cdot k^{n-1}\ll p\ll p^{\frac{n-1}{2}}$
when $n\geq 3$.
Finally, when $n=2$, let us deal with $t=1$ without loss of generality. In this situation, one has $m_{1},x_{1}\not\equiv 0\,(\textsf{mod }p)$. Note that
$j\equiv a_{1}x_{1}^{k}+a_{2}x_{2}^{k}\equiv a_{1}x_{1}^{k}+a_{2}(m_{1}a_{2})^{-1}m_{2}a_{1}x_{1}^{k-1}x_{2}\equiv a_{1}x_{1}^{k-1}\big{(}x_{1}+m_{1}^{-1}m_{2}x_{2}\big{)}\,(\textsf{mod }p).$
If $m_{2}\equiv 0\,(\textsf{mod }p)$, then the above equivalence has at most $k-1$ solutions in $x_{1}\,(\textsf{mod }p)$. Otherwise we have $x_{2}\equiv m_{1}m_{2}^{-1}(ja_{1}^{-1}x_{1}^{-(k-1)}-x_{1})\,(\textsf{mod }p)$. Then
$(m_{1}a_{2})^{-1}m_{2}a_{1}x_{1}^{k-1}\equiv x_{2}^{k-1}\equiv\left(m_{1}m_{2}^{-1}(ja_{1}^{-1}x_{1}^{-(k-1)}-x_{1})\right)^{k-1}\,(\textsf{mod }p),$
i.e.,
$\big{(}-(m_{1}a_{2})^{-1}m_{2}a_{1}+(-1)^{k-1}\big{)}x_{1}^{k(k-1)}+(-1)^{k-2}m_{1}^{k-1}m_{2}^{-(k-1)}(k-1)(ja_{1}^{-1})x_{1}^{k(k-2)}$
$\qquad+\sum\limits_{i=2}^{k-1}(-1)^{k-1-i}{k-1\choose i}(ja_{1}^{-1})^{i}x_{1}^{k(k-1-i)}\equiv 0\,(\textsf{mod }p).$
This involves a non-zero polynomial of degree no larger than $k(k-1)$. So there are at most $k(k-1)$ solutions in $x_{1}\,(\textsf{mod }p)$. Now we conclude that $\mathfrak{N}\leq 2k(k-1)^{2}\ll 1$.
Corollary 1.5 then follows from Theorem 1.4. ∎
Proof of Theorem 1.6.
For sufficiently large $p$, we have
$(\nabla F)(x)=(k_{1}a_{1}x_{1}^{k_{1}-1},\ldots,k_{n}a_{n}x_{n}^{k_{n}-1})\not\equiv\mathbf{0}\,(\textsf{mod }p)$
when $\mathbf{x}\not\equiv\mathbf{0}\,(\textsf{mod }p)$. Hensel’s lemma can be applied on the roots of $F(\mathbf{x})\equiv j\,(\textsf{mod }p^{l})$ $(l\geq 1)$. Combining Lemma 2.3, one can obtain that $|S_{r,j}|\ll q^{n-1}$. Then we deduce similarly as in the proof of Theorem 1.4 that
$\mathcal{N}_{r,j}\ll_{F}q^{-1}|E_{1}||E_{2}|+q^{2n}\cdot\sum\limits_{\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\widehat{1_{E_{2}}}(\mathbf{m})\widehat{1_{S_{r,j}}}(\mathbf{m}).$
Now
$\sum\limits_{j\,(\textsf{mod }q)}|\mathcal{N}_{r,j}|^{2}$ $\ll\Sigma_{1}+\Sigma_{2},$
where
$\Sigma_{1}=\sum\limits_{j\,(\textsf{mod }q)}\left(q^{-1}|E_{1}||E_{2}|\right)^{2}=q^{-1}|E_{1}|^{2}|E_{2}|^{2},$
and
$\Sigma_{2}=q^{4n}\sum\limits_{\mathbf{m},\mathbf{l}\not\equiv\mathbf{0}\,(\textsf{mod }q)}\overline{\widehat{1_{E_{1}}}(\mathbf{m})}\widehat{1_{E_{2}}}(\mathbf{m})\widehat{1_{E_{1}}}(\mathbf{l})\overline{\widehat{1_{E_{2}}}(\mathbf{l})}\sum\limits_{j\,(\textsf{mod }q)}\widehat{1_{S_{r,j}}}(\mathbf{m})\overline{\widehat{1_{S_{r,j}}}(\mathbf{l})}.$
Note that
$\Sigma_{3}:=$ $\sum\limits_{j\,(\textsf{mod }q)}\widehat{1_{S_{r,j}}}(\mathbf{m})\overline{\widehat{1_{S_{r,j}}}(\mathbf{l})}$
$=\frac{1}{q^{2n}}\sum\limits_{j\,(\textsf{mod }q)}\sum\limits_{\mathbf{x}\,(\textsf{mod }q)\atop F(\mathbf{x})\equiv j\,(\textsf{mod }q)}\sum\limits_{\mathbf{y}\,(\textsf{mod }q)\atop F(\mathbf{y})\equiv j\,(\textsf{mod}q)}e_{q}(-\mathbf{m}\cdot\mathbf{x}+\mathbf{l}\cdot\mathbf{y})$
$=\frac{1}{q^{2n}}\sum\limits_{\mathbf{x},\mathbf{y}\,(\textsf{mod}q)\atop F(\mathbf{x})\equiv F(\mathbf{y})\,(\textsf{mod }q)}e_{q}(-\mathbf{m}\cdot\mathbf{x}+\mathbf{l}\cdot\mathbf{y})$
$=\frac{1}{q^{2n+1}}\sum\limits_{s\,(\textsf{mod }q)}\sum\limits_{\mathbf{x},\mathbf{y}\,(\textsf{mod }q)}e_{q}(-\mathbf{m}\cdot\mathbf{x}+\mathbf{l}\cdot\mathbf{y}+sF(\mathbf{x})-sF(\mathbf{y}))$
$=\frac{1}{q^{2n+1}}\sum\limits_{s\not\equiv 0\,(\textsf{mod }q)}\sum\limits_{\mathbf{x},\mathbf{y}\,(\textsf{mod }q)}e_{q}(-\mathbf{m}\cdot\mathbf{x}+\mathbf{l}\cdot\mathbf{y}+sF(\mathbf{x})-sF(\mathbf{y})),$
where the last two steps follows from the orthogonality and the fact that $\mathbf{m},\mathbf{l}\not\equiv\mathbf{0}\,(\textsf{mod }q)$.
Noting that the inner summation is a product of exponential sums in one variable, one has
$|\Sigma_{3}|\leq\frac{1}{q^{2n+1}}\cdot\sum\limits_{s\not\equiv 0\,(\textsf{mod }q)}\left|\prod\limits_{i=1}^{n}\sum\limits_{x_{i}\,(\textsf{mod }q)}e_{q}\big{(}-m_{i}x_{i}+sa_{i}x_{i}^{k_{i}}\big{)}\cdot\prod\limits_{i=1}^{n}\sum\limits_{y_{i}\,(\textsf{mod }q)}e_{q}\big{(}l_{i}y_{i}-sa_{i}y_{i}^{k_{i}}\big{)}\right|.$
For $(m_{i},s),(l_{i},s)\not\equiv(0,0)\,(\textsf{mod }p)$, the condition (5) in Lemma 2.6 follows from Lemma 2.2, and the condition (6) in Lemma 2.6 is confirmed by observing that
#$\#\{x_{i}(\textsf{mod }p):\,-m_{i}+sa_{i}k_{i}x_{i}^{k_{i}-1}\equiv 0\,(\textsf{mod }p)\},\,\#\{y_{i}(\textsf{mod }p):\,-l_{i}+sa_{i}k_{i}y_{i}^{k_{i}-1}\equiv 0\,(\textsf{mod }p)\}\leq k_{i}-1.$
For $s\not\equiv 0\,(\textsf{mod }q)$, we deduce from Lemma 2.6 that
$\left|\sum\limits_{x_{i}\,(\textsf{mod }q)}e_{q}\big{(}-m_{i}x_{i}+sa_{i}x_{i}^{k_{i}}\big{)}\right|,\,\,\left|\sum\limits_{y_{i}\,(\textsf{mod }q)}e_{q}\big{(}l_{i}y_{i}-sa_{i}y_{i}^{k_{i}}\big{)}\right|\,\,\ll_{k}\,qp^{-1/2}$ (10)
for any $m_{i},l_{i}\,(\textsf{mod }q)$ with $1\leq i\leq n$. Moreover, let us assume without loss of generality that $k_{1}=k_{\ast}=\min\nolimits_{1\leq i\leq n}k_{i}$. Then
$|\Sigma_{3}|\ll q^{-2n-1}\cdot\left(qp^{-1/2}\right)^{2n-2}\cdot\sum\limits_{s\not\equiv 0\,(\textsf{mod }q)}\left|\sum\limits_{x_{1}\,(\textsf{mod }q)}e_{q}\big{(}-m_{1}x_{1}+sa_{1}x_{1}^{k_{1}}\big{)}\cdot\sum\limits_{y_{1}\,(\textsf{mod }q)}e_{q}\big{(}l_{1}y_{1}-sa_{1}y_{1}^{k_{1}}\big{)}\right|$
$\leq q^{-3}p^{-n+1}\cdot\left(\sum\limits_{s\,(\textsf{mod }q)}\left|\sum\limits_{x_{1}\,(\textsf{mod }q)}e_{q}\big{(}-m_{1}x_{i}+sa_{1}x_{1}^{k_{1}}\big{)}\right|^{2}\right)^{\frac{1}{2}}\left(\sum\limits_{s\,(\textsf{mod }q)}\left|\sum\limits_{y_{1}\,(\textsf{mod }q)}e_{q}\big{(}-l_{1}x_{1}+sa_{1}y_{1}^{k_{1}}\big{)}\right|^{2}\right)^{\frac{1}{2}}$
$\leq q^{-3}p^{-n+1}\cdot q^{3-2/k_{\ast}}=q^{-2/k_{\ast}}p^{-n+1},$ (11)
where we have applied (10), the Cauchy-Schwartz inequality and Lemma 2.7.
It follows that
$\Sigma_{2}$ $\ll q^{4n}\cdot q^{-2/k_{\ast}}p^{-n+1}\cdot\prod\limits_{i=1,2}\left(\sum\limits_{\mathbf{m}(\textsf{mod }q)}|\widehat{1_{E_{i}}}(\mathbf{m})|^{2}\right)^{1/2}\cdot\prod\limits_{i=1,2}\left(\sum\limits_{\mathbf{l}(\textsf{mod }q)}|\widehat{1_{E_{i}}}(\mathbf{l})|^{2}\right)^{1/2}$
$\ll q^{4n}\cdot q^{-2/k_{\ast}}p^{-n+1}\cdot q^{-2n}|E_{1}||E_{2}|=q^{2n-2/k\ast}p^{-n+1}|E_{1}||E_{2}|.$
As a result, we have
$\sum\limits_{j\,(\textsf{mod }q)}\mathcal{N}_{r,j}^{2}\ll q^{-1}|E_{1}|^{2}|E_{2}|^{2}+q^{2n-2/k_{\ast}}p^{-n+1}|E_{1}||E_{2}|.$
Moreover, since
$|E_{1}|^{2}|E_{2}|^{2}=\left(\sum\limits_{j\in\Delta_{n,r}(E_{1},E_{2})}\mathcal{N}_{r,j}\right)^{2}\leq|\Delta_{n,r}(E_{1},E_{2})|\cdot\sum\limits_{j\,(\textsf{mod }q)}\mathcal{N}_{r,j}^{2},$
we conclude that
$|\Delta_{n,r}(E_{1},E_{2})|\geq\frac{|E_{1}|^{2}|E_{2}|^{2}}{\sum\limits_{j\,(\textsf{mod }q)}\mathcal{N}_{r,j}^{2}}\gg\min\{q,\,q^{-2n+2/k_{\ast}}p^{n-1}|E_{1}||E_{2}|\}.$
In particular, one has $|\Delta_{n,r}(E_{1},E_{2})|\gg q$ if
$\frac{\sqrt{|E_{1}||E_{2}|}}{q^{n}}\gg q^{\frac{1}{2}-\frac{1}{k_{\ast}}}p^{-\frac{n-1}{2}}.$
This completes the proof. ∎
Remark 3.1.
For the case $r=1$, we may use (7) instead of (8) in the above proof, and obtain $\ll p^{-n}$ at the end of (11). As a consequence, the same result (1) as in [11] can be obtained.
4 Extension estimates associated to circles and orbits
Let $V$ be a variety over $\mathbb{F}_{p}^{n}$. The $L^{u}\to L^{u^{\prime}}$ Fourier extension problem for $V$ asks us to determine all exponents $1\leq u,u^{\prime}\leq\infty$ such that the following inequality
$||(fd\sigma)^{\vee}||_{L^{u^{\prime}}(\mathbb{F}_{p}^{n},dm)}\leq C||f||_{L^{u}(V,d\sigma)},$ (12)
holds for some constant $C>0$ and all complex valued functions $f$ on $V$. Here $dm$ is the counting measure on $\mathbb{F}_{p}^{n}$, $d\sigma$ is the normalized surface measure on $V$, and
$(fd\sigma)^{\vee}(m)=\frac{1}{|V|}\sum_{x\in V}f(x)e_{p}(m\cdot x),$
$||f||_{L^{u}(V,d\sigma)}=\left(\frac{1}{|V|}\sum_{x\in V}|f(x)|^{u}\right)^{1/u},$
and
$||g||_{L^{u^{\prime}}(\mathbb{F}_{p}^{n},dm)}=\left(\sum_{m\in\mathbb{F}_{p}^{n}}|g(m)|^{u^{\prime}}\right)^{1/u^{\prime}},$
for any functions $f$ and $g$. The $L^{u}\to L^{u^{\prime}}$ Fourier extension problem has been studied intensively in the literature for paraboloids, spheres, cones, and homogeneous varieties with applications in several areas of Mathematics including Discrete Geometry and Combinatorial Number Theory. We refer the interested reader to a series of papers [8, 9, 10, 12, 13, 14] for more discussions.
When $V$ is a circle in the plane $\mathbb{F}_{p}^{2}$, it has been proved in [2] that
$||(fd\sigma)^{\vee}||_{L^{4}(\mathbb{F}_{p}^{2},dm)}\ll||f||_{L^{2}(V,d\sigma)}.$
In this paper, to improve Corollary 1.5 in two dimensions, we need to study the finite p-adic ring analog of this estimate. The extension estimates are of independent interest and are expected to have more applications in other topics.
4.1 Extension theorems
In the rest of this section, we always assume that $n=2$. Recall that
$C_{r,j}=\{\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}:\,\|\mathbf{x}\|\equiv j\,(\textsf{mod }p^{r})\}.$
We denote by $d\sigma_{r,j}$ the normalised surface measure on $C_{r,j}$.
Theorem 4.1.
Let $r\geq 1$ be an integer. Let $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. Then
$\Big{(}\sum_{m\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}}|(fd\sigma_{r,j})^{\vee}(m)|^{4}\Big{)}^{1/2}\ll p^{-\frac{r+1}{2}}\sum_{x\in C_{r,j}}|f(x)|^{2}.$
We now define a more general extension problem. Denote
$G_{r}:=SO_{2}(\mathbb{Z}/p^{r}\mathbb{Z})=\left\{\begin{bmatrix}a&-b\\ b&a\end{bmatrix}\,(\textsf{mod }p^{r}):\,\,a^{2}+b^{2}\equiv 1\,(\textsf{mod }p^{r})\right\}$
for $r\geq 1$. The group $G_{r}$ acts on $(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ naturally by $\mathbf{x}\mapsto\theta\mathbf{x}$, where $\theta\in G_{r}$ and $\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$. Write the orbit of $\mathbf{x}$ by
$\textsf{orb}_{r}(\mathbf{x})=\big{\{}\theta\mathbf{x}:\,\theta\in G_{r}\big{\}},$
and the stabilizer of $\mathbf{x}$ is
$\textsf{stab}_{r}(\mathbf{x})=\{\theta\in G_{r}:\,\theta\mathbf{x}\equiv\mathbf{x}\,(\textsf{mod }p^{r})\}.$
For $\mathbf{m}=(m_{1},m_{2})\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, recall that $v_{\mathbf{m}}=\min\{\textsf{ord}_{p}(m_{1}),\textsf{ord}_{p}(m_{2})\}$.
Theorem 4.2.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$ be an integer. Let $\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ be such that $\mathbf{m}\neq\mathbf{0}$. Denote by $d{\color[rgb]{0,0,0}\sigma_{r}}$ the normalized surface measure on $\textsf{orb}_{r}(\mathbf{m})$. Then
$\Big{(}\sum_{x\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}}|(fd\sigma_{r})^{\vee}(x)|^{4}\Big{)}^{1/2}\ll p^{-\frac{r-3v_{\mathbf{m}}+1}{2}}\sum_{x\in\textsf{orb}_{r}(\mathbf{m})}|f(x)|^{2}.$
Theorem 4.3.
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$ be an integer. Let $\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ be such that $\mathbf{m}\neq\mathbf{0}$. Denote by $d\sigma_{r}$ the normalized surface measure on $\textsf{orb}_{r}(\mathbf{m})$. Write $\mathbf{m}=p^{v_{m}}\tilde{\mathbf{m}}$ with $\tilde{\mathbf{m}}\in(\mathbb{Z}/p^{r-v_{m}}\mathbb{Z})^{2}$ and $\|\tilde{\mathbf{m}}\|\not\equiv 0\,(\textsf{mod }p)$. Then
$\Big{(}\sum_{x\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}}|(fd\sigma_{r})^{\vee}(x)|^{4}\Big{)}^{1/2}\ll p^{-\frac{\color[rgb]{0,0,0}r-3v_{\mathbf{m}}+1}{2}}\sum_{x\in\textsf{orb}_{r}(\mathbf{m})}|f(x)|^{2}.$
4.2 Preliminary lemmas
To prove these extension theorems, we first need to collect and prove a number of preliminary results. All detailed proofs in this subsection are written under the assumption that $p\equiv 3\,(\textsf{mod }4)$. In the case $p\equiv 1\,(\textsf{mod }4)$, the arguments are identical but more complicated, and will be omitted in this paper.
Let $p$ be a prime with $p\equiv 3\,(\textsf{mod }4)$, $r\geq 1$ be an integer. For any $l\in\mathbb{Z}$, one has $(l,p)=1$ if and only if $(l,p^{r})=1$. For any $\theta\in G_{r}$ and $\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, it is not hard to verify that $\|\theta\mathbf{x}\|\equiv\|\mathbf{x}\|\quad(\textsf{mod }p^{r})$. Moreover, we have $v_{\theta\mathbf{x}}\geq v_{\mathbf{x}}$. Since $\theta$ is invertible, one obtains that $v_{\theta\mathbf{x}}=v_{\mathbf{x}}$.
It is well-known (see [5] for example) that
$|C_{1,j}|=|G_{1}|=p+1,\quad(j\in(\mathbb{Z}/p\mathbb{Z})^{\ast}),\qquad|C_{1,0}|=1$
[…]
Lemma 4.4.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$. Then $|G_{r}|=p^{r}(1+1/p)$.
Proof.
For $r=1$, we have $|G_{1}|=p+1$. Now consider the circumstances that $r\geq 2$. For any $\theta=\begin{bmatrix}a&-b\\ b&a\end{bmatrix}\in G_{r}$, there is some $\theta_{0}=\begin{bmatrix}a_{0}&-b_{0}\\ b_{0}&a_{0}\end{bmatrix}\in G_{1}$ such that $\theta\equiv\theta_{0}\,(\textsf{mod }p)$. Applying Lemma 2.1 to $(a_{0},b_{0})$ and the polynomial $F(x,y)=x^{2}+y^{2}-1$. Then $(\nabla F)(a_{0},b_{0})=(2a_{0},2b_{0})\not\equiv\mathbf{0}\,(\textsf{mod }p)$, since $a_{0}^{2}+b_{0}^{2}\equiv 1\,(\textsf{mod }p)$. Thus,
#$\#\left\{(z_{1},z_{2})\,(\textsf{mod }p^{r-1}):\,(a_{0}+pz_{1})^{2}+(b_{0}+pz_{2})^{2}\equiv 1\,(\textsf{mod }p^{r})\right\}=p^{r-1}.$
It follows that
$|G_{r}|=p^{r-1}|G_{1}|=p^{r}(1+1/p).$
Lemma 4.5.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$. For any $\mathbf{0}\neq\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, let $\mathbf{x}=p^{v_{\mathbf{x}}}\tilde{\mathbf{x}}$ for some $\tilde{\mathbf{x}}\in(\mathbb{Z}/p^{r-v_{\mathbf{x}}}\mathbb{Z})^{2}$ with $v_{\tilde{\mathbf{x}}}=0$. Then
$\textsf{orb}_{r}(\mathbf{x})=\big{\{}p^{v_{\mathbf{x}}}\theta_{0}\tilde{\mathbf{x}}:\,\theta_{0}\in G_{r-v_{\mathbf{x}}}\big{\}}=p^{v_{\mathbf{x}}}C_{r-v_{\mathbf{x}},\|\tilde{\mathbf{x}}\|}.$
And
$|\textsf{stab}_{r}(\mathbf{x})|=p^{v_{\mathbf{x}}},\qquad|\textsf{orb}_{r}(\mathbf{x})|=p^{r-v_{\mathbf{x}}}(1+1/p).$
Proof.
Write $v=v_{\mathbf{x}}$ for simplicity. Here $10\leq v\leq r-$ 1. Let $\tilde{\mathbf{x}}=(\tilde{x_{1}},\tilde{x_{2}})$. The equation $\theta\mathbf{x}\equiv\mathbf{x}\,(\textsf{mod }p^{r})$ is equivalent to
$\begin{bmatrix}a-1&-b\\ b&a-1\end{bmatrix}\begin{bmatrix}p^{v}\tilde{x_{1}}\\ p^{v}\tilde{x_{2}}\end{bmatrix}\equiv\begin{bmatrix}0\\ 0\end{bmatrix}\quad(\textsf{mod }p^{r}),$ (13)
or equivalently,
$\begin{bmatrix}\tilde{x_{1}}&-\tilde{x_{2}}\\ \tilde{x_{2}}&\tilde{x_{1}}\end{bmatrix}\begin{bmatrix}a-1\\ b\end{bmatrix}\equiv\begin{bmatrix}0\\ 0\end{bmatrix}\quad(\textsf{mod }p^{r-v}).$ (14)
Since $v_{\tilde{\mathbf{x}}}=0$, one has $\tilde{x_{1}}^{2}+\tilde{x_{2}}^{2}\not\equiv 0\,(\textsf{mod }p)$ (recalling that $p\equiv 3\,(\textsf{mod }4)$). The coefficient matrix is invertible and
$\begin{bmatrix}a\\ b\end{bmatrix}\equiv\begin{bmatrix}1\\ 0\end{bmatrix},\quad(\textsf{mod }p^{r-v}).$
By Lemma 2.1, the number of $(a,b)\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ satisfying the above equivalence and $a^{2}+b^{2}\equiv 1\,(\textsf{mod }p^{r})$ is exactly $p^{v}$. So
$\big{|}\textsf{stab}_{r}(\mathbf{x})\big{|}=p^{v}.$
It then follows that $|\textsf{orb}_{r}(\mathbf{x})|=|G_{r}|/|\textsf{stab}_{r}(\mathbf{x})|=p^{r-v}(1+1/p)$.
Moreover, for any $\theta\in G_{r}$, there is some $\theta_{0}\in G_{r-v}$ such that $\theta\equiv\theta_{0}\,(\textsf{mod }p^{r-v})$. It can be verified that $\theta\tilde{\mathbf{x}}\equiv\theta_{0}\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})$. So
$\textsf{orb}_{r}(\mathbf{x})=\big{\{}\theta(p^{v}\tilde{\mathbf{x}}):\,\theta\in G_{r}\big{\}}=\big{\{}p^{v}\theta_{0}\tilde{\mathbf{x}}:\,\theta_{0}\in G_{r-v}\big{\}}.$
Note that $|G_{r-v}|=p^{r-v}(1+1/p)$ by Lemma 4.4, the elements on the right-hand side of above formula give different members of the orbit. Furthermore, one has $\|\tilde{\mathbf{x}}\|\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}$, since $v_{\tilde{\mathbf{x}}}=0$ and $p\equiv 3\,(\textsf{mod }4)$. We have $\{\theta_{0}\tilde{\mathbf{x}}:\,\theta_{0}\in G_{r-v}\}=C_{r-v,\|\tilde{\mathbf{x}}\|}$. The proof is completed. ∎
Lemma 4.6.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$. Let $j$ be an integer with $\text{ord}_{p}(j)=v$. Then
$|C_{r,j}|=\begin{cases}p^{r}(1+1/p),\quad&\text{if }0\leq v<r\text{ and }v\text{ is even},\\ p^{2\lfloor r/2\rfloor},\quad&\text{if }v=r,\\ 0,\quad&\text{otherwise}.\end{cases}$
More concretely, we have
$C_{r,j}=\begin{cases}\left\{p^{u}\mathbf{z}+p^{r-u}\mathbf{w}:\,\mathbf{z}\in C_{r-2u,\tilde{j}},\,\mathbf{w}\in\mathbb{Z}/p^{u}\mathbb{Z}\right\},\quad&\text{if }j=p^{2u}\tilde{j},\,0\leq u\leq\lfloor\frac{r-1}{2}\rfloor,\,\tilde{j}\in(\mathbb{Z}/p^{r-2u}\mathbb{Z})^{\ast}\\ \left\{p^{\lceil r/2\rceil}\mathbf{w}:\,\mathbf{w}\in\mathbb{Z}/p^{\lfloor r/2\rfloor}\mathbb{Z}\right\},&\text{if }j=0,\\ \emptyset,&\text{otherwise}.\end{cases}$
Proof.
We first consider the case $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. By applying Lemma 2.1, similar arguments as previous show that
$|C_{r,j}|=p^{r-1}|C_{1,j}|=p^{r}(1+1/p).$
Next, consider the circle $C_{r,0}$ $(r\geq 2)$. Since $C_{1,0}=\{\mathbf{0}\}$, we can write
$C_{r,0}=\{p\mathbf{y}:\,\mathbf{y}\in C^{\prime}_{r,0}\},\quad C^{\prime}_{r,0}=\big{\{}\mathbf{y}\,(\textsf{mod }p^{r-1}):\,\|p\mathbf{y}\|\equiv 0\,(\textsf{mod }p^{r})\big{\}}.$
When $r=2$, one sees that $C_{2,0}=\{p\mathbf{y}:\,\mathbf{y}\in(\mathbb{Z}/pZ)^{2}\}$ and $|C_{2,0}|=p^{2}$. When $r\geq 3$, we further write $\mathbf{y}=\mathbf{z}+p^{r-2}\mathbf{w}$ with $\mathbf{z}\in(\mathbb{Z}/p^{r-2}\mathbb{Z})^{2}$ and $\mathbf{w}\in(\mathbb{Z}/p\mathbb{Z})^{2}$. The condition $\|p\mathbf{y}\|\equiv\mathbf{0}\,(\textsf{mod }p^{r})$ holds if and only if $\|\mathbf{z}+p^{r-2}\mathbf{w}\|=\|\mathbf{y}\|\equiv\mathbf{0}\,(\textsf{mod }p^{r-2})$, if and only if $\|\mathbf{z}\|\equiv\mathbf{0}\,(\textsf{mod }p^{r-2})$. It follows that
$C_{r,0}=\big{\{}p(\mathbf{z}+p^{r-2}\mathbf{w}):\,\mathbf{z}\in C_{r-2,0},\,\mathbf{w}\in(\mathbb{Z}/p\mathbb{Z})^{2}\big{\}},$
and $|C_{r,0}|=p^{2}|C_{r-2,0}|$.
By induction, we conclude that
$C_{r,0}=\left\{p^{\lceil r/2\rceil}\mathbf{w}:\,\mathbf{w}\in\mathbb{Z}/p^{\lfloor r/2\rfloor}\mathbb{Z}\right\}$
and $|C_{r,0}|=p^{2\lfloor r/2\rfloor}$.
[…]
$C_{r,j}=\{p\mathbf{y}:\,\mathbf{y}\in C^{\prime}_{r,j}\},\quad C^{\prime}_{r,j}=\big{\{}\mathbf{y}\,(\textsf{mod }p^{r-1}):\,\|p\mathbf{y}\|\equiv p^{v}\tilde{j}\,(\textsf{mod }p^{r})\big{\}}.$
When $v=1$, it is easy to see that $C^{\prime}_{r,j}=\emptyset$. When $v\geq 2$, which means $r\geq 3$, similar arguments as previous show that
$C_{r,p^{v}\tilde{j}}=\big{\{}p(\mathbf{z}+p^{r-2}\mathbf{w}):\,\mathbf{z}\in C_{r-2,p^{v-2}\tilde{j}},\,\mathbf{w}\in(\mathbb{Z}/p\mathbb{Z})^{2}\big{\}},$
and $|C_{r,p^{v}\tilde{j}}|=p^{2}\cdot|C_{r-2,p^{v-2}\tilde{j}}|$. By induction, it follows that
$|C_{r,p^{v}\tilde{j}}|=\begin{cases}p^{r}(1+1/p),\quad&\text{if }v\text{ is even},\\ 0,\quad&\text{if }v\text{ is odd}\end{cases}$
when $11\leq v\leq r-$ 1. Indeed,
$C_{r,p^{2u}\tilde{j}}=\left\{p^{u}\mathbf{z}+p^{r-u}\mathbf{w}:\,\mathbf{z}\in C_{r-2u,\tilde{j}},\,\mathbf{w}\in(\mathbb{Z}/p^{u}\mathbb{Z})^{2}\right\}$ (15)
when $r\geq 3$ and $21\leq u\leq(r-1)/$ 2. ∎
Remark:
When $p\equiv 3\,(\textsf{mod }4)$, we have the disjoint union
$(\mathbb{Z}/p^{r}\mathbb{Z})^{2}=\left(\bigcup\limits_{u=0}^{\lfloor(r-1)/2\rfloor}\bigcup\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-2u}\mathbb{Z})^{\ast}}C_{r,p^{2u}\tilde{j}}\right)\,\cup\,C_{r,0},$
whose cardinalities give
$p^{2r}$ $=\sum\limits_{u=0}^{\lfloor(r-1)/2\rfloor}p^{r}(1+1/p)\cdot p^{r-2u}(1-1/p)+p^{2\lfloor r/2\rfloor}$
$=\sum\limits_{u=0}^{\lfloor(r-1)/2\rfloor}(p^{2r-2u}-p^{2r-2u-2})+p^{2\lfloor r/2\rfloor}.$
Lemma 4.7.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$. Let $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. For any $\mathbf{0}\neq\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, we have
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in C_{r,j}^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}\,\leq\,2p^{r-1}.$
Proof.
The set on the left-hand side involves system of congruences
$\begin{cases}x_{1}^{2}+x_{2}^{2}\equiv j,\\ y_{1}^{2}+y_{2}^{2}\equiv j,\\ x_{1}-y_{1}\equiv z_{1},\\ x_{2}-y_{2}\equiv z_{2}.\end{cases}$ (16)
Considering the system (16) modulo $p$, one sees that
$2y_{1}z_{1}+2y_{2}z_{2}+z_{1}^{2}+z_{2}^{2}\equiv 0\,(\textsf{mod }p).$
First, we consider the case $\mathbf{z}\not\equiv\mathbf{0}\,(\textsf{mod }p)$. Assume without loss of generality that $z_{1}\not\equiv 0\,(\textsf{mod }p)$. Inserting $y_{1}\equiv-(2z_{1})^{-1}(2y_{2}z_{2}+z_{1}^{2}+z_{2}^{2})\,(\textsf{mod }p)$ into $y_{1}^{2}+y_{2}^{2}\equiv j\,(\textsf{mod }p)$, one obtains
$z_{1}^{-2}(z_{1}^{2}+z_{2}^{2})y_{2}^{2}+z_{1}^{-2}z_{2}(z_{1}^{2}+z_{2}^{2})y_{2}+4^{-1}z_{1}^{-2}(z_{1}^{2}+z_{2}^{2})^{2}\equiv j\quad(\textsf{mod }p).$
When $p\equiv 3\,(\textsf{mod }4)$, $\mathbf{z}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ implies that $z_{1}^{2}+z_{2}^{2}\not\equiv 0\,(\textsf{mod }p)$. The left-hand side of above congruence has degree $2$. So there are at most two solutions in $y_{2}$. Moreover, each choice of $y_{2}$ exactly determines the choices of $x_{1}$, $y_{1}$ and $x_{2}$.
Now we apply Hensel’s lemma. The Jacobian matrix, in $(x_{1},y_{1},x_{2},y_{2})$, is given by
$\begin{bmatrix}2x_{1}&0&2x_{2}&0\\ 0&2y_{1}&0&2y_{2}\\ 1&-1&0&0\\ 0&0&1&-1\end{bmatrix}.$
By elementary operations, we obtain
$\begin{bmatrix}1&0&-1&0\\ 0&1&0&-1\\ 0&0&x_{1}&x_{2}\\ 0&0&y_{1}&y_{2}\end{bmatrix}.$
Since $j\not\equiv 0\,(\textsf{mod }p)$, it satisfies that $\mathbf{x},\mathbf{y}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ and the rank modulo $p$ is at least $3$. Then the conclusion follows from Lemma 2.1.
[…]
$(y_{1}+p^{k}\tilde{z_{1}})^{2}+(y_{2}+p^{k}\tilde{z_{2}})^{2}\equiv j\equiv y_{1}^{2}+y_{2}^{2},\,(\textsf{mod }p^{k+1}),$
i.e., $2y_{1}\tilde{z_{1}}+2y_{2}\tilde{z_{2}}\equiv 0\,(\textsf{mod }p)$. Noting that $v_{\tilde{\mathbf{z}}}=0$, we assume without loss of generality that $\tilde{z_{1}}\not\equiv 0\,(\textsf{mod }p)$. Then
$x_{1}\equiv y_{1}\equiv y_{2}\tilde{z_{2}}\tilde{z_{1}}^{-1}\equiv x_{2}\tilde{z_{2}}\tilde{z_{1}}^{-1},\quad(\textsf{mod }p).$
Now any solution $(\mathbf{x},\mathbf{y})$ to the system (16) modulo $p^{r}$ satisfies that
$\begin{cases}x_{1}^{2}+x_{2}^{2}\equiv j,\\ x_{1}\equiv x_{2}\tilde{z_{2}}\tilde{z_{1}}^{-1},\\ x_{1}\equiv y_{1},\\ x_{2}\equiv y_{2},\end{cases}\quad(\textsf{mod }p).$
It is not hard to see that there are at most $2$ solutions modulo $p$. Now we lift these solutions modulo $p$ to solutions to (16) modulo $p^{r}$. By applying Hensel’s lemma with the rank of Jacobian matrix being $3$, we conclude that
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in C_{r,j}^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}\leq 2p^{r-1}.$
Lemma 4.8.
Let $p\equiv 3\,(\textsf{mod }4)$ and $r\geq 1$. Let $\mathbf{m},\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ be such that $\mathbf{m},\mathbf{z}\neq\mathbf{0}$. Then
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in(\textsf{orb}_{r}(\mathbf{m}))^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}\,\leq\,2p^{r-v_{\mathbf{m}}-1}.$
Proof.
Write $v=v_{\mathbf{m}}$ and $\mathbf{m}=p^{v}\tilde{\mathbf{m}}$. Here $10\leq v\leq r-$ 1, $\tilde{\mathbf{m}}\in(\mathbb{Z}/p^{r-v_{\mathbf{m}}}\mathbb{Z})^{2}$ and $v_{\tilde{\mathbf{m}}}=0$. Write $\|\tilde{\mathbf{m}}\|=\tilde{j}$ for simplicity. Then $\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}$. By Lemma 4.5, we have $\textsf{orb}(\mathbf{m})=p^{v}C_{r-v,\tilde{j}}$. Write $\mathbf{x}=p^{v}\tilde{\mathbf{x}}$ and $\mathbf{y}=p^{v}\tilde{\mathbf{y}}$ with $\tilde{\mathbf{x}},\tilde{\mathbf{y}}\in C_{r-v,\tilde{j}}$. When $\mathbf{z}\not\equiv\mathbf{0}(\textsf{mod }p^{v})$, it is obvious that
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in(\textsf{orb}_{r}(\mathbf{m}))^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}=0.$
[…]
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in(\textsf{orb}_{r}(\mathbf{m}))^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}=\#\big{\{}(\tilde{\mathbf{x}},\tilde{\mathbf{y}})\in(C_{\tilde{j},r-v})^{2}:\,\tilde{\mathbf{x}}-\tilde{\mathbf{y}}\equiv\tilde{\mathbf{z}}\,(\textsf{mod }p^{r-v})\big{\}}.$
By Lemma 4.7, the above quantity can be bounded by $2p^{r-v-1}$. ∎
With the same argument, we obtain similar results for the case $p\equiv 1\,(\textsf{mod }4)$.
Lemma 4.9.
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$. For any $0\neq\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, let $\mathbf{x}=p^{v_{\mathbf{x}}}\tilde{\mathbf{x}}$ for some $\tilde{\mathbf{x}}\in(\mathbb{Z}/p^{r-v_{\mathbf{x}}}\mathbb{Z})^{2}$ with $v_{\tilde{\mathbf{x}}}=0$. Then
$|\textsf{stab}_{r}(\mathbf{x})|=p^{v_{\mathbf{x}}},\qquad|\textsf{orb}_{r}(\mathbf{x})|=p^{r-v_{\mathbf{x}}}(1-1/p).$
And $\textsf{orb}_{r}(\mathbf{x})=p^{v_{\mathbf{x}}}\textsf{orb}_{r-v_{\mathbf{x}}}(\tilde{\mathbf{x}})$.
Lemma 4.10.
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$. Let $j$ be an integer with $\text{ord}_{p}(j)=v$. Then
$|C_{r,j}|=\begin{cases}(v+1)p^{r}(1-1/p),\quad&\text{if }0\leq v<r,\\ (rp+p-r)p^{r-1},\quad&\text{if $v=r$}\end{cases}$
Lemma 4.11.
Let $p\equiv 1\,(\textsf{mod }4)$ and $r\geq 1$. Let $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}$. For any $\mathbf{0}\neq\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, we have
#$\#\big{\{}(\mathbf{x},\mathbf{y})\in C_{r,j}^{2}:\,\mathbf{x}-\mathbf{y}\equiv\mathbf{z}\,(\textsf{mod }p^{r})\big{\}}\,\leq\,2p^{r-1}.$
Lemma 4.12.
For $p\equiv 1\,(\textsf{mod }4)$ be a prime, $r\geq 1$ be an integer. Let $\mathbf{m}=p^{v_{\mathbf{m}}}\tilde{\mathbf{m}}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\|\tilde{\mathbf{m}}\|\not\equiv 0\,(\textsf{mod }p)$. For any $\mathbf{z}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, we have
#$\#\{(\mathbf{x},\mathbf{y})\in(\textsf{orb}_{r}(\mathbf{m}))^{2}\colon\mathbf{x}-\mathbf{y}\equiv\mathbf{z}(\textsf{mod }p^{r})\}\ll p^{r-v_{\mathbf{m}}-1}.$
4.3 Proof of extension theorems
Theorems 4.1 and 4.2 are proved in the same way. Let $V$ be a variety with $V\subseteq(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, which can be $C_{r,j}$ or $\textsf{orb}_{r}(\mathbf{m})$, on which the normalized surface measure is denoted by $d\sigma$. For simplicity, we denote $q=p^{r}$ in this subsection.
We have
$\sum_{\mathbf{m}\,(\textsf{mod }q)}|(fd\sigma)^{\vee}(\mathbf{m})|^{4}=\sum_{\mathbf{m}\,(\textsf{mod }q)}\left|\frac{1}{|V|}\sum_{\mathbf{x}\in V}f(\mathbf{x})e_{q}(\mathbf{m}\cdot\mathbf{x})\right|^{4}$
$\qquad=\frac{q^{2}}{|V|^{4}}\sum_{\xi,\xi^{\prime},\eta,\eta^{\prime}\in V\atop\xi-\eta\equiv\xi^{\prime}-\eta^{\prime}\,(\textsf{mod }q)}f(\xi)f(\xi^{\prime})\overline{f(\eta)f(\eta^{\prime})}=\frac{q^{2}}{|V|^{4}}\sum_{\zeta\,(\textsf{mod }q)}\left|\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}f(\xi)\overline{f(\eta)}\right|^{2}.$
For $\zeta\equiv\mathbf{0}\,(\textsf{mod }q)$, we have
$\Big{|}\sum_{\xi,\eta\in V\atop\xi\equiv\eta\,(\textsf{mod }q)}f(\xi)\overline{f(\eta)}\Big{|}^{2}=\Big{(}\sum_{\xi\in V}|f(\xi)|^{2}\Big{)}^{2}.$
For $\zeta\not\equiv\mathbf{0}\,(\textsf{mod }p^{r})$, the Cauchy-Schwarz inequality implies
$\sum_{\zeta\not\equiv\mathbf{0}\,(\textsf{mod }q)}\Big{|}\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}f(\xi)\overline{f(\eta)}\Big{|}^{2}\leq\sum_{\zeta\not\equiv\mathbf{0}\,(\textsf{mod }q)}\Big{(}\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}1^{2}\Big{)}\Big{(}\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}|f(\xi)|^{2}|f(\eta)|^{2}\Big{)}.$
Assuming that $\sum\nolimits_{\xi,\eta\in V\,\xi-\eta\equiv\zeta\,(\textsf{mod }q)}1\ll U$ for all $\zeta\not\equiv\mathbf{0}\,(\textsf{mod }q)$, then
$\sum_{\zeta\not\equiv\mathbf{0}\,(\textsf{mod }q)}\Big{|}\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}f(\xi)\overline{f(\eta)}\Big{|}^{2}\ll U\,\sum_{\zeta\,(\textsf{mod }p^{r})}\Big{(}\sum_{\xi,\eta\in V\atop\xi-\eta\equiv\zeta\,(\textsf{mod }q)}|f(\xi)|^{2}|f(\eta)|^{2}\Big{)}=U\,\Big{(}\sum_{\xi\in V}|f(\xi)|^{2}\Big{)}^{2}.$
It follows that
$\left(\sum_{m\,(\textsf{mod }q)}|(fd\sigma_{r,j})^{\vee}(m)|^{4}\right)^{1/2}\ll\frac{q{\color[rgb]{0,0,0}(1+U)}^{1/2}}{|V|^{2}}\sum_{x\in V}|f(x)|^{2}.$
On the one hand, to get the bound $U$, we use Lemmas 4.7 and 4.8. On the other hand, Lemma 4.6 gives estimates on the size of $V$. Hence, the theorems follow.
5 Proof of Theorems 1.8 and 1.9
These two theorems are proved by the same argument.
By Lemma 4.6, it is clear that that the number of pairs $(\mathbf{x},\mathbf{y})\in E\times E$ such that $||\mathbf{x}-\mathbf{y}||\not\in(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$ is much smaller than $|E|^{2}$, since $|E|\gg p^{2r-1}$. By the Cauchy-Schwarz inequality, we have
$|E|^{4}$ $\ll\Big{(}\sum\limits_{\mathbf{x},\mathbf{y}\in E\atop\|\mathbf{x}-\mathbf{y}\|\in(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}}1\Big{)}^{2}=\big{(}\sum\limits_{j\in\Delta_{2,r}(E)\cap(\mathbb{Z}/p^{r}\mathbb{Z})^{\ast}}\sum\limits_{\mathbf{x}\in E}\sum\limits_{\mathbf{y}\in E\atop\|\mathbf{x}-\mathbf{y}\|\equiv j\,(\textsf{mod }p^{r})}1\Big{)}^{2}$
$\leq|\Delta_{2,r}(E)|\cdot|E|\cdot\sum\limits_{j\not\equiv 0\,(\textsf{mod }p)}\sum\limits_{\mathbf{x}\in E}\Big{(}\sum\limits_{\mathbf{y}\in E\atop\|\mathbf{x}-\mathbf{y}\|\equiv j\,(\textsf{mod }p^{r})}1\Big{)}^{2}.$
So
$|\Delta_{2,r}(E)|\gg\frac{|E|^{3}}{\mathcal{N}},$
where
$\mathcal{N}:=\sum\limits_{j\not\equiv 0\,(\textsf{mod }p)}\sum\limits_{\mathbf{x}\in E}\Big{(}\sum\limits_{\mathbf{y}\in E\atop\|\mathbf{x}-\mathbf{y}\|\equiv j\,(\textsf{mod }p^{r})}1\Big{)}^{2}=\sum\limits_{x\in E}\,\sum\limits_{j\not\equiv 0\,(\textsf{mod }p)}\,\Big{(}\sum_{\mathbf{z}\in C_{r,j}}1_{E}(\mathbf{x}-\mathbf{z})\Big{)}^{2}.$
For any given $\mathbf{x}\in E$, we have
$\sum_{j\not\equiv 0\,(\textsf{mod }p)}\Big{(}\sum_{\mathbf{z}\in C_{r,j}}1_{E}(\mathbf{x}-\mathbf{z})\Big{)}^{2}=\sum_{j\not\equiv 0\,(\textsf{mod }p)}\sum_{\mathbf{y},\mathbf{z}\in C_{r,j}}1_{E}(\mathbf{x}-\mathbf{y})1_{E}(\mathbf{x}-\mathbf{z})$
$\leq\sum_{j\not\equiv 0\,(\textsf{mod }p)}\sum_{\mathbf{z}\in C_{r,j}}1_{E}(\mathbf{x}-\mathbf{z})\sum_{\theta\in G_{r}}1_{E}(\mathbf{x}-\theta\mathbf{z})\leq\sum_{{\color[rgb]{0,0,0}\mathbf{z}\,(\textsf{mod }p^{r})}}\sum_{\theta\in G_{r}}1_{E}(\mathbf{x}-\mathbf{z})1_{E}(\mathbf{x}-\theta\mathbf{z}).$
In the next step, we write
$\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}\sum_{\theta\in G_{r}}1_{E}(\mathbf{x}-\mathbf{z})1_{E}(\mathbf{x}-\theta\mathbf{z})$
$\quad=\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}\sum_{\theta\in G_{r}}\sum\limits_{\mathbf{m}\,(\textsf{mod }p^{r})}\widehat{1_{E}}(\mathbf{m})e_{p^{r}}(\mathbf{m}\cdot(\mathbf{x}-\mathbf{z}))\overline{\sum\limits_{\mathbf{m}^{\prime}\,(\textsf{mod }p^{r})}\widehat{1_{E}}(\mathbf{m}^{\prime})e_{p^{r}}(\mathbf{m}^{\prime}\cdot(\mathbf{x}-\theta\mathbf{z}))}$
$\quad=\sum_{\theta\in G_{r}}\sum\limits_{\mathbf{m}\,(\textsf{mod}p^{r})}\sum\limits_{\mathbf{m}^{\prime}\,(\textsf{mod }p^{r})}\widehat{1_{E}}(\mathbf{m})\overline{\widehat{1_{E}}(\mathbf{m}^{\prime})}e_{p^{r}}\big{(}(\mathbf{m}-\mathbf{m}^{\prime})\cdot\mathbf{x}\big{)}\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}e_{p^{r}}\big{(}-(\mathbf{m}\cdot\mathbf{z}-\mathbf{m}^{\prime}\cdot(\theta\mathbf{z})\big{)}.$
Note that
$\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}e_{p^{r}}\big{(}-(\mathbf{m}\cdot\mathbf{z}-\mathbf{m}^{\prime}\cdot(\theta\mathbf{z})\big{)}=\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}e_{p^{r}}(-(\mathbf{m}-\theta^{-1}\mathbf{m}^{\prime})\cdot\mathbf{z}\big{)}=\begin{cases}p^{2r},\quad&\text{if }\mathbf{m}^{\prime}=\theta\mathbf{m},\\ 0,&\text{otherwise}.\end{cases}$
One deduces that
$\mathcal{N}\ll$ $\sum\limits_{x\in E}\sum_{\mathbf{z}\,(\textsf{mod }p^{r})}\sum_{\theta\in G_{r}}1_{E}(\mathbf{x}-\mathbf{z})1_{E}(\mathbf{x}-\theta\mathbf{z})$
$=$ $p^{2r}\sum\limits_{x\in E}\sum_{\mathbf{m}\,(\textsf{mod }p^{r})}\sum_{\theta\in G_{r}}\widehat{1_{E}}(\mathbf{m})\overline{\widehat{1_{E}}(\theta\mathbf{m})}e_{p^{r}}((\mathbf{m}-\theta\mathbf{m})\cdot\mathbf{x})$
$\ll$ $\sum\limits_{x\in E}\left(\frac{|E|^{2}}{p^{r}}+\mathcal{E}(\mathbf{x})\right)=\frac{|E|^{3}}{p^{r}}+\sum\limits_{x\in E}\mathcal{E}(\mathbf{x}),$
where
$\mathcal{E}(\mathbf{x}):=p^{2r}\sum\limits_{v=0}^{r-1}\mathcal{E}_{v}(\mathbf{x}),\qquad\mathcal{E}_{v}(\mathbf{x})=\sum_{\mathbf{m}\,(\textsf{mod }p^{r})\atop v_{\mathbf{m}}=v}\sum_{\theta\in G_{r}}\widehat{1_{E}}(\mathbf{m})\overline{\widehat{1_{E}}(\theta\mathbf{m})}e_{p^{r}}((\mathbf{m}-\theta\mathbf{m})\cdot\mathbf{x}).$
[…]
Write $\mathbf{m}=p^{v}\tilde{\mathbf{m}}$ and $\tilde{\theta}\equiv\theta\,(\textsf{mod }p^{r-v})$. Then $\tilde{\theta}\tilde{\mathbf{m}}\equiv\theta{\color[rgb]{0,0,0}\tilde{\mathbf{m}}}\,(\textsf{mod }p^{r-v})$. Noting that $|G_{r}/G_{r-v}|=p^{v}$, we have
$\mathcal{E}_{v}(\mathbf{x})$ $=p^{v}\sum\limits_{\tilde{\mathbf{m}}\,(\textsf{mod }p^{r-v})\atop v_{\tilde{\mathbf{m}}=0}}\sum_{\tilde{\theta}\in G_{r-v}}\widehat{1_{E}}(p^{v}\tilde{\mathbf{m}})\overline{\widehat{1_{E}}(p^{v}\tilde{\theta}\tilde{\mathbf{m}})}e_{p^{r-v}}\big{(}(\tilde{\mathbf{m}}-\tilde{\theta}\tilde{\mathbf{m}})\cdot\mathbf{x})\big{)}$ (17)
$=p^{v}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}\,}\sum\limits_{\tilde{\mathbf{m}}\in C_{r-v,\tilde{j}}}\sum\limits_{\mathbf{m}\in C_{r-v,\tilde{j}}}\widehat{1_{E}}(p^{v}\tilde{\mathbf{m}})\overline{\widehat{1_{E}}(p^{v}\mathbf{m})}e_{p^{r-v}}\big{(}(\tilde{\mathbf{m}}-\mathbf{m})\cdot\mathbf{x})\big{)}$
$=p^{v}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}\,}\left|\sum\limits_{\tilde{m}\in C_{r-v,\tilde{j}}}\widehat{1_{E}}(p^{v}\tilde{\mathbf{m}})e_{p^{r-v}}(\tilde{\mathbf{m}}\cdot\mathbf{x})\right|^{2}.$
When $11\leq v\leq r-$ 1, we rewrite
$\widehat{1_{E}}(p^{v}\tilde{\mathbf{m}})$ $=\frac{1}{p^{2r}}\sum\limits_{\mathbf{y}\,(\textsf{mod }p^{r})}1_{E}(\mathbf{y})e_{p^{r}}(-p^{v}\tilde{\mathbf{m}}\cdot\mathbf{y})$
$={\color[rgb]{0,0,0}\frac{1}{p^{2(r-v)}}}\sum\limits_{\mathbf{y}_{1}\,(\textsf{mod }p^{r-v})}{\color[rgb]{0,0,0}\frac{1}{p^{2v}}}\sum\limits_{\mathbf{y}_{2}\,(\textsf{mod }p^{v})}1_{E}(\mathbf{y}_{1}+p^{r-v}\mathbf{y}_{2})e_{p^{r-v}}(-\tilde{\mathbf{m}}\cdot\mathbf{y}_{1}):=\widehat{g_{E,r-v}}(\tilde{\mathbf{m}}).$
Here, the Fourier transformation on the left-hand side is over $(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$, while that on the right-hand side is over $(\mathbb{Z}/p^{r-v}\mathbb{Z})^{2}$. And, for $\mathbf{y}_{1}\,(\textsf{mod }p^{r-v})$,
$g_{E,r-v}(\mathbf{y}_{1})=\frac{1}{p^{2v}}\sum\limits_{\mathbf{y}_{2}\,(\textsf{mod }p^{v})}1_{E}(\mathbf{y}_{1}+p^{r-v}\mathbf{y}_{2})=\frac{1}{p^{2v}}\#\{\mathbf{y}\in E:\,\mathbf{y}\equiv\mathbf{y}_{1}\,(\textsf{mod }p^{r-v})\}.$
We note that for all $\mathbf{y}_{1}\,(\textsf{mod }p^{r-v})$, one has $g_{E,r-v}(\mathbf{y}_{1})\leq 1$. When $v=0$, the above notation gives $g_{E,r}=1_{E}$. Now , for all $10\leq v\leq r-$ 1, we have
$\mathcal{E}_{v}(\mathbf{x})=p^{v}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}\,}\left|\sum\limits_{\tilde{m}\in C_{r-v,\tilde{j}}}\widehat{g_{E,r-v}}(\tilde{\mathbf{m}})e_{p^{r-v}}(\tilde{\mathbf{m}}\cdot\mathbf{x})\right|^{2}$
Write $f=\widehat{g_{E,r-v}}\big{|}_{C_{r-v,\tilde{j}}}$. Denote
$(fd\sigma_{r-v,\tilde{j}})^{\vee}(x)=\frac{1}{|C_{r-v,\tilde{j}}|}\sum\limits_{\mathbf{y}\in C_{r-v,\tilde{j}}}f(y)e_{p^{r-v}}(\mathbf{y}\cdot\mathbf{x}).$
It follows that
$\mathcal{E}_{v}(\mathbf{x})\ll p^{2r-v}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}\,}\left|(fd\sigma_{\color[rgb]{0,0,0}r-v,\tilde{j}})^{\vee}(\mathbf{x})\right|^{2}.$
[…]
$\color[rgb]{0,0,0}E(\textsf{mod }p^{\gamma}):=\{\mathbf{x}\,(\textsf{mod }p^{\gamma}):\,\mathbf{x}\in E\}=\{\tilde{\mathbf{x}}\,(\textsf{mod }p^{\gamma}):\,\tilde{\mathbf{x}}\equiv\mathbf{x}\,(\textsf{mod }p^{\gamma})\text{ for some}\mathbf{x}\in E\}$
for $1\leq\gamma\leq r$. Then, by Cauchy-Schwartz inequality, Theorem 4.1 and Parseval’s identity, one deduces that
$p^{2r}\sum\limits_{\mathbf{x}\in E}\mathcal{E}_{v}(\mathbf{x})$ $\ll p^{4r-v}\sum\limits_{\tilde{\mathbf{x}}\in E(\textsf{mod }p^{r-v})}p^{2v}g_{E,r-v}(\tilde{\mathbf{x}})\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}}|(fd\sigma_{r-v,\tilde{j}})^{\vee}(\tilde{\mathbf{x}})|^{2}$
$\ll p^{4r+v}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}}\left(\sum\limits_{\tilde{\mathbf{x}}\in E(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{x})|^{2}\right)^{1/2}\left(\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})}|(fd\sigma_{r-v,\tilde{j}})^{\vee}(\tilde{\mathbf{x}})|^{4}\right)^{1/2}$
$\ll p^{4r+v}\left(\sum\limits_{\tilde{\mathbf{x}}\in E(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{\mathbf{x}})|^{2}\right)^{1/2}\cdot p^{-\frac{r-v+1}{2}}\sum\limits_{\tilde{j}\in(\mathbb{Z}/p^{r-v}\mathbb{Z})^{\ast}}\sum\limits_{\tilde{\mathbf{y}}\in C_{r-v,\tilde{j}}}|\widehat{g_{E,r-v}}(\tilde{\mathbf{y}})|^{2}$
${\color[rgb]{0,0,0}\ll p^{\frac{7r+3v-1}{2}}\left(\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{\mathbf{x}})|^{2}\right)^{1/2}\cdot\sum\limits_{\tilde{\mathbf{y}}\,(\textsf{mod }p^{r-v})}|\widehat{g_{E,r-v}}(\tilde{\mathbf{y}})|^{2}}$
$\ll p^{\frac{3r+7v-1}{2}}\left(\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{\mathbf{x}})|^{2}\right)^{3/2}.$
The last sum can be bounded trivially by
$\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{\mathbf{x}})|^{2}\leq\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p^{r-v})}|g_{E,r-v}(\tilde{\mathbf{x}})|\leq\frac{|E|}{p^{2v}}.$
In other words,
$p^{2r}\sum\limits_{\mathbf{x}\in E}\mathcal{E}_{v}(\mathbf{x})\ll p^{\frac{3r+v-1}{2}}|E|^{3/2}.$
With this bound in hand, one has
$\mathcal{N}\ll\frac{|E|^{3}}{p^{r}}+p^{\frac{3r-1}{2}}|E|^{3/2}\sum_{v=0}^{r-1}p^{v/2}\ll\frac{|E|^{3}}{p^{r}}+p^{2r-1}|E|^{3/2}\ll\frac{|E|^{3}}{p^{r}},$
whenever $|E|\gg p^{2r-\frac{2}{3}}$. This proves Theorem 1.8.
Indeed,
$p^{2r}\sum\limits_{\mathbf{x}\in E}\sum\limits_{v=0}^{r-2}\mathcal{E}_{v}(\mathbf{x})\ll p^{\frac{3r-1}{2}}|E|^{3/2}\sum_{v=0}^{r-2}p^{v/2}\ll p^{2r-\frac{3}{2}}|E|^{3/2}\ll\frac{|E|^{3}}{p^{r}}$
whenever $|E|\gg p^{2r-1}$. In the statement of Theorem 1.9, we know that
#$\#\{(\mathbf{x}_{1},\mathbf{x}_{2})\in E^{2}:\,\mathbf{x}_{1}\equiv\mathbf{x}_{2}\,(\textsf{mod }p)\}\ll p^{2r-\frac{7}{3}}|E|.$
Then
$\sum\limits_{\tilde{\mathbf{x}}\,(\textsf{mod }p)}|g_{E,1}(\tilde{\mathbf{x}})|^{2}=\frac{1}{p^{4r-4}}\#\{(\mathbf{x}_{1},\mathbf{x}_{2})\in E^{2}:\,\mathbf{x}_{1}\equiv\mathbf{x}_{2}\,(\textsf{mod }p)\}\ll p^{-2r+\frac{5}{3}}|E|.$
It follows that
$p^{2r}\sum\limits_{x\in E}\mathcal{E}_{r-1}(\mathbf{x})\ll p^{5r-4}\cdot\left(p^{-2r+\frac{5}{3}}|E|\right)^{3/2}\ll p^{2r-\frac{3}{2}}|E|^{\frac{3}{2}}\ll\frac{|E|^{3}}{p^{r}},$
and then $\mathcal{N}\ll\frac{|E|^{3}}{p^{r}}$. Plugging this bound to the above argument gives us Theorem 1.9.
Proof of Corollary 1.10.
Under the condition of Corollary 1.10, we have
#$\#\{(\mathbf{x}_{1},\mathbf{x}_{2})\in E^{2}:\,\mathbf{x}_{1}\equiv\mathbf{x}_{2}\,(\textsf{mod }p)\}=\sum\limits_{\mathbf{x}_{1}\in E}\#\{\mathbf{x}_{2}\in E:\,\mathbf{x}_{2}\equiv\mathbf{x}_{1}\,(\textsf{mod }p)\}\ll p^{2r-7/3}|E|.$
Then the conclusion follows from Theorem 1.9. ∎
6 Proof of Theorems 1.11, 1.12, 1.13, and 1.14
In this section, we write $q=p^{r}$ for simplicity. To prove Theorems 1.11 and 1.13, we use a general framework developed for pseudo-random graphs in which Lemma 2.5 plays a crucial role.
Proof of Theorem 1.11:
Let $G=(V,E)$ be a graph with the vertex set $V$ and the edge set $E$. Let $M$ be its adjacency matrix, i.e., $M_{ij}=1$ if and only if there is an edge between $i$ and $j$ and zero otherwise. Let $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{|V|}$ be its eigenvalues. Assume that $|V|=N$, the graph $G$ is called a $(N,d,\lambda)$-graph if each vertex is of degree $d$ and the second eigenvalue, determined by $\max\{\lambda_{2},-\lambda_{N}\}$, is at most $\lambda$.
Let $G$ and $H$ be two graphs. The Cartesian product of $G$ and $H$, denoted by $G\square H$, is defined as follows. The vertex set $V(G\square H)=V(G)\times V(H)$, and there is an edge between $(u_{1},v_{1})$ and $(u_{2},v_{2})$ if and only if either $u_{1}=u_{2}$ and $(v_{1},v_{2})\in E(H)$ or $v_{1}=v_{2}$ and $(u_{1},u_{2})\in V(G)$.
We recall the following result on the number of rectangles in a Cartesian product of graphs in [21].
Theorem 6.1.
Let $G_{i}$ be $(N_{i},d_{i},\lambda_{i})$-graphs with $1\leq i\leq 2$. Set $G=G_{1}\square G_{2}$. For any $10<\delta^{\prime}<\delta<$ 1, there exists $\epsilon>0$ such that for any $A\subset V(G_{1}\square G_{2})$ with $|A|\geq\delta|V(G_{1}\square G_{2})|$, if $\max\left\{\frac{\lambda_{1}}{d_{1}},\frac{\lambda_{2}}{d_{2}}\right\}<\epsilon$, then
$N=\sum_{(u_{1},u_{2})\in E(G_{1}),(v_{1},v_{2})\in E(G_{2})}A(u_{1},v_{1})A(u_{1},v_{2})A(u_{2},v_{1})A(u_{2},v_{2})>\delta^{\prime 4}N_{1}N_{2}d_{1}d_{2}.$
Given $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$. To prove Theorem 1.11, we will apply the above theorem with appropriate graphs $G$ and $H$. More precisely, define $G=H$ with $V(G)=V(H)=(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, and there is an edge between two vertices $\mathbf{x}$ and $\mathbf{y}$ if and only if $F(\mathbf{x}-\mathbf{y})\equiv j\,(\textsf{mod }q)$. So $G$ and $H$ have $q^{n}$ vertices and are regular of degree $(1+o(1))q^{n-1}$. On the other hand, since $G$ and $H$ are Cayley graphs with the generating set $S=\{\mathbf{x}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\colon F(\mathbf{x})\equiv j\,(\textsf{mod }q)\}$. Thus, all non-trivial eigenvalues of $G$ and $H$ are bounded by $q^{n}\cdot\max_{\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }q)}|\widehat{1_{S}}(\mathbf{m)}|\ll q^{n-1}p^{-\frac{n-1}{2}}$ by Lemma 2.5. In other words, $G$ and $H$ are $\big{(}q^{n},(1+o(1))q^{n-1},q^{n-1}p^{-\frac{n-1}{2}}\big{)}$-graphs.
Thus, Theorem 1.11 follows directly from Theorem 6.1 with $A=E$.
Proof of Theorem 1.13:
To prove Theorem 1.13, we follow the same approach as in the proof of Theorem 1.11. More precisely, we make use of the following result on $(N,d,\lambda)$-graphs, which is also taken from [21].
Theorem 6.2.
Let $G$ be a $(N,d,\lambda)$-graph, $\ell\geq 1$ an integer, and $A$ be a vertex set with $|A|\gg_{\ell}\lambda\cdot\frac{N}{d}$. Let $P_{\ell}(U)$ denotes the number of paths of length $\ell$ in $A$. Then we have
$P_{\ell}(E)\sim\frac{|A|^{\ell+1}d^{\ell}}{N^{\ell}}.$
With the graph $G$ defined as above, Theorem 1.13 follows from Theorem 6.2 immediately.
Proof of Theorem 1.12:
We first recall the statement of Theorem 1.4: for $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ if $|E|\gg q^{n}p^{-\frac{n-1}{2}}$, then for any $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$, the number of pairs $(\mathbf{x},\mathbf{y})\in E\times E$ such that $F(\mathbf{x}-\mathbf{y})\equiv j\,(\textsf{mod }q)$ is at least $\gg q^{-1}|E|^{2}$.
By the pigeon-hole principle, there exists $\mathbf{u}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ such that $F(\mathbf{u})\equiv j\,(\textsf{mod }q)$ and the number of pairs $(\mathbf{x},\mathbf{y})\in E\times E$ such that $\mathbf{x}=\mathbf{y}+\mathbf{u}$ is at least $\gg\frac{|E|^{2}}{q^{n}}$. Let $E_{1}\subset E$ be the set of $\mathbf{y}\in E$ such that $\mathbf{y}+\mathbf{u}\in E$. Without loss of generality, we assume that $|E_{1}|=|E|^{2}/q^{n}$. By Theorem 1.4 again, the number of pairs $(\mathbf{y}_{1},\mathbf{y}_{2})\in E_{1}\times E_{2}$ such that $F(\mathbf{y}_{1}-\mathbf{y}_{2})\equiv j\,(\textsf{mod }q)$ is at least $|E_{1}|^{2}/q=|E|^{4}/q^{2n+1}$. Among these pairs, the number of pairs with $\mathbf{y}_{1}=\mathbf{y}_{2}+\mathbf{u}$ is at most $|E_{1}|$, which is smaller than $|E_{1}|^{2}/q$. Therefore, by the pigeon-hole principle again, there exists $\mathbf{u}^{\prime}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ such that $\mathbf{u}\neq\mathbf{u}^{\prime}$, $F(\mathbf{u}^{\prime})\equiv j\,(\textsf{mod }q)$, and the number of pairs $(\mathbf{y}_{1},\mathbf{y}_{2})\in E_{1}\times E_{1}$ such that $\mathbf{y}_{1}=\mathbf{y}_{2}+\mathbf{u}^{\prime}$ is at least $\gg\frac{|E|^{2}}{q^{n}}>0$.
With each pair $(\mathbf{y}_{1},\mathbf{y}_{2})$ satisfying the above property, we have a cycle of the from $(\mathbf{y}_{1},\mathbf{y}_{2},\mathbf{y}_{2}+\mathbf{u},\mathbf{y}_{1}+\mathbf{u})$. This completes the proof.
Proof of Theorem 1.14:
To prove Theorem 1.14, we need a stronger version of Theorem 1.4, namely, the pinned distance version.
Theorem 6.3.
Let $E_{1},E_{2}\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $|E_{1}|,|E_{2}|\gg q^{n}p^{-\frac{n-1}{2}}$, then there exists $E_{1}^{\prime}\subset E_{1}$ with $|E_{1}^{\prime}|\gg|E_{1}|$ and $|\Delta_{n.r,\mathbf{x}}(E_{2})|\gg q$ for all $\mathbf{x}\in E_{1}^{\prime}$. Here
$\Delta_{n.r,\mathbf{x}}(E_{2}):=\{F(\mathbf{x}-\mathbf{y})\colon\mathbf{y}\in E_{2}\}.$
To prove this strong version, we introduce an incidence problem between points and $F$-spheres in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$.
A $F$-sphere centered at $x$ of radius $j$ is the set
$\{y\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\colon F(x-y)\equiv j\textsf{mod }q\}.$
Let $\mathcal{P}$ be a set of points in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ and $\mathcal{S}$ be a set of $F$-spheres. The number of incidences between $\mathcal{P}$ and $\mathcal{S}$, denoted by $I(\mathcal{P},\mathcal{S})$, is defined by
$I(\mathcal{P},\mathcal{S})=\#\{(p,s)\in\mathcal{P}\times\mathcal{S}\colon p\in s\}.$
Theorem 6.4.
Let $\mathcal{P}$ be a set of points in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ and $\mathcal{S}$ be a set of $F$-spheres with radii in $(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$. Then the number of incidences between $\mathcal{P}$ and $\mathcal{S}$ satisfies
$I(\mathcal{P},\mathcal{S})\leq\frac{|\mathcal{P}||\mathcal{S}|}{q}+Cq^{n-\frac{1}{2}}p^{-\frac{n-1}{2}}|\mathcal{P}|^{1/2}|cS|^{1/2}.$
Proof.
We partition $\mathcal{S}$ into $\mathcal{S}_{i}$, $i\in(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$, such that spheres in $\mathcal{S}_{i}$ are of radius $i$. By abuse of notation, we denote the set of centers in $\mathcal{S}_{i}$ by $\mathcal{S}_{i}$. By Theorem 1.4, we have
$I(\mathcal{P},\mathcal{S}_{i})\leq\frac{|\mathcal{P}||\mathcal{S}_{i}|}{q}+q^{n-1}p^{-\frac{n-1}{2}}|\mathcal{P}|^{1/2}|\mathcal{S}_{i}|^{1/2}.$
Taking the sum over all $i$ and by using the Cauchy-Schwarz inequality, we have the conclusion. ∎
Proof of Theorem 6.3.
For $\mathbf{x}\in E_{1}$, let $\Delta_{n,r,{\mathbf{x}}}^{*}(E_{2})$ be the set of distances in $(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$ between $\mathbf{x}$ and points in $E_{2}$. Let $\mathcal{S}(\mathbf{x})$ be the set of spheres centered at $\mathbf{x}$ of radii in $\Delta_{n.r,\mathbf{x}}^{*}(E_{2})$, and set $\mathcal{S}=\cup_{\mathbf{x}\in E_{1}}\mathcal{S}(\mathbf{x})$. We have $|\mathcal{S}|=\sum_{\mathbf{x}\in E_{1}}|\Delta_{n,r,{\mathbf{x}}}^{*}(E_{2})|$. Under the conditions on the sizes of $E_{1}$ and $E_{2}$, for any $j\in(\mathbb{Z}/p^{r}\mathbb{Z})^{*}$, it follows from Theorem 1.4 that the number of pairs $(\mathbf{x},\mathbf{y})\in E_{1}\times E_{2}$ such that $F(\mathbf{x}-\mathbf{y})\equiv j\,(\textsf{mod }q)$ is at least $|E_{1}||E_{2}|/2q$.
Thus, $I(E_{2},\mathcal{S})\geq|E_{1}||E_{2}|/4$. Applying Theorem 6.4 gives
$I(E_{2},\mathcal{S})\leq\frac{|E_{2}|\sum_{\mathbf{x}\in E_{1}}|\Delta_{\mathbf{x}}(E_{2})^{*}|}{q}+Cq^{n-\frac{1}{2}}p^{-\frac{n-1}{2}}|E_{2}|^{1/2}\left(\sum_{\mathbf{x}\in E_{1}}|\Delta_{n.r,\mathbf{x}}(E_{2})^{*}|\right)^{1/2}.$
If $\sum_{\mathbf{x}\in E_{1}}|\Delta_{n.r,\mathbf{x}}(E_{2})^{*}|\leq|E_{1}|q/8$, then $|E_{1}||E_{2}|\leq 8C^{2}q^{2n}p^{-(n-1)}$, so we reach a contradiction as long as $|E_{1}||E_{2}|>8C^{2}q^{2n}p^{-(n-1)}$.
Thus, if $|E_{1}||E_{2}|>8C^{2}q^{2n}p^{-(n-1)}$, then we have
$\sum_{\mathbf{x}\in E_{1}}|\Delta_{n.r,\mathbf{x}}(E_{2})^{*}|>\frac{|E_{1}|q}{8}.$ (18)
Set $E_{1}^{\prime}:=\{\mathbf{x}\in E_{1}\colon|\Delta_{\mathbf{x}}(E_{2})^{*}|\geq q/32\}$. We are going to show that $|E_{1}^{\prime}|\geq|E_{1}|/32$. Indeed, otherwise, we have
$\sum_{\mathbf{x}\in E_{1}}|\Delta_{\mathbf{x}}(E_{2})^{*}|=\sum_{\mathbf{x}\in E_{1}^{\prime}}+\sum_{\mathbf{x}\in E_{1}\setminus E_{1}^{\prime}}\leq q|E_{1}^{\prime}|+\frac{q|E_{1}|}{32}\leq\frac{q|E_{1}|}{16},$
which contradicts (18). Therefore, we have proved that there exists $E_{1}^{\prime}\subset E_{1}$ with $|E_{1}^{\prime}|\geq|E_{1}|/32$ and for all $\mathbf{x}\in E_{1}^{\prime}$, we have $|\Delta_{n.r,\mathbf{x}}(E_{2})^{*}|\geq q/32$. The proof is complete. ∎
With Theorem 6.3 in hand, Theorem 1.14 now follows from a combinatorial argument which is identical to [22, Proof of Theorem 3.1] with Theorem 6.3 in place of [22, Lemma 2.2]. So we omit the details.
7 Sharpness examples
In this section, we provide some sharpness examples for results stated in the introduction.
Example 7.1.
Let $p$ be a large prime. Suppose that $n$ is odd and $-1$ is a $k$-th power modulo $p$, i.e., $\xi^{k}\equiv-1\,(\textsf{mod }p)$. We consider the polynomial $F(\mathbf{x})=x_{1}^{k}+x_{2}^{k}+\ldots+x_{n}^{k}$.
For $r=1$, let us take
$\Pi_{0}=\{(u,\xi u):\,u\in\mathbb{Z}/p\mathbb{Z}\}.$
Then, for any $(u,\xi u)\in\Pi_{0}$, one has $\|(u,\xi u)\|=u^{k}+(\xi u)^{k}\equiv 0\,(\textsf{mod }p)$. This is a one-dimensional subspace of $(\mathbb{Z}/p\mathbb{Z})^{2}$. We have $\|(u,\xi u)-(u^{\prime},\xi u^{\prime})\|=0$ for any $(u,\xi u)$, $(u^{\prime},\xi u^{\prime})\in\Pi_{0}$. Let $21\leq l<p/$ 2 be a parameter to be determined later. Take
$E_{1}=\big{\{}(x_{1},x_{2},\ldots,x_{n})\in\mathbb{Z}/p\mathbb{Z}:\,x_{1}\in\{1,2,\ldots,l\},\,(x_{2},x_{3}),\ldots,(x_{n-1},x_{n})\in\Pi_{0}\big{\}}.$
Then $|E_{1}|=lp^{\frac{n-1}{2}}$, and
$|\Delta_{n,1}(E_{1})|=\#\left\{(-l+1)^{k},(-l+2)^{k},\ldots,(l-1)^{k}\right\}<2l.$
For $r\geq 2$, let us take
$E_{r}=\big{\{}\mathbf{x}+p\mathbf{y}:\,\mathbf{x}\in E_{1},\,\mathbf{y}\in(\mathbb{Z}/p^{r-1}\mathbb{Z})^{n}\big{\}}$
as a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$. It is not hard to see that $|E_{r}|=lp^{nr-\frac{n+1}{2}}$. Note that, for any two elements $\mathbf{z}=\mathbf{x}+p\mathbf{y},\,\mathbf{z}^{\prime}=\mathbf{x}^{\prime}+p\mathbf{y}^{\prime}\in E_{r}$, we have $\|\mathbf{z}-\mathbf{z}^{\prime}\|\equiv\|\mathbf{x}-\mathbf{x}^{\prime}\|\,(\textsf{mod }p)$. So
$|\Delta_{n,r}(E_{r})|\leq p^{r-1}\cdot|\Delta_{n,1}(E_{1})|<2lp^{r-1}.$
To sum up, if we take $l=\lceil p\cdot w(p)\rceil$ with $w(p)=\textit{o}(1)$ $(p\rightarrow\infty)$ being a positive-valued function tending to 00 arbitrarily slowly, then there are sets $E_{r}$ with density $p^{-\frac{n-1}{2}}w(p)$ such that $|\Delta_{n,r}(E_{r})|<2p^{r}w(p)=\textit{o}(p^{r})$. So Corollary 1.5 is optimal in general.
Example 7.2.
Given a positive integer $C$, we first claim that there exists a set $E_{1}\subset(\mathbb{Z}/p\mathbb{Z})^{n}$ with $\delta_{E_{1}}\geq C^{-1}p^{-n/2}$ such that $\Delta_{n,1}(E_{1})=\{0\}$ or $|\Delta_{n,1}(E_{1})|\leq(p+1)/C$.
If $-1$ is a square or $n\equiv 0\,(\textsf{mod }4)$, as above, we can find such a set $E_{1}\subset(\mathbb{Z}/p\mathbb{Z})^{n}$ of size $p^{n/2}$ such that $\Delta_{n,1}(E_{1})=\{0\}$.
If $-1$ is a non-square and $n\equiv 2\,(\textsf{mod }4)$, then let $\theta$ be a generator of the group $G_{1}=SO_{2}(\mathbb{F}_{p})$, which is cyclic of order $p+1$, and let $v_{0}\in(\mathbb{Z}/p\mathbb{Z})^{2}$ such that $||v_{0}||=1$. By the Dirichlet’s theorem on arithmetic progressions, one can choose $p$ large enough such that $C$ divides $p+1$. Let $G$ be the subgroup spanned by $\theta^{C}$. Then we have $|G|=(p+1)/C$. Set $X=\textsf{orb}_{1}(\mathbf{v}_{0})$. A direct computation shows that $X$ satisfies the desired property, namely, $|\Delta_{2,1}(X)|\leq|X|=(p+1)/C$. Let $X^{\prime}\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n-2}$ of size $p^{(n-2)/2}$ such that $\Delta_{n,1}(X^{\prime})=\{0\}$. Set $E_{1}=X\times X^{\prime}$. Then we have $|E_{1}|=p^{\frac{n-2}{2}}\cdot\frac{p+1}{C}$. By the definition of $E_{1}$, we obtain $|\Delta_{n,1}(E_{1})|\leq(p+1)/C$.
For $r\geq 2$, we set
$E_{r}=\{\mathbf{x}+((\mathbb{Z}/p^{r}\mathbb{Z})\setminus U_{r})^{n}\colon\mathbf{x}\in E_{1}\}.$
So
$\delta_{E_{r}}=p^{-rn}|E|\geq p^{-rn}p^{\frac{n-2}{2}}\cdot\frac{p+1}{C}\cdot p^{(r-1)n}\geq C^{-1}p^{-n/2},$
and
$|\Delta_{n,r}(E_{r})|\leq p^{r-1}|\Delta_{n,1}(E_{1})|\ll p^{r}/C.$
8 Acknowledgements
Thang Pham would like to thank the Alfréd Rényi Institute of Mathematics for the hospitality and excellent working conditions. Thang Pham was partially supported by ERC grant “GeoScape”, no. 882971, under Prof. János Pach. We thank Le Quang Hung for checking all proofs in Section 4 for the case $p\equiv 1\,(\textsf{mod }4)$.